### v1

# Polarons and Exciton-Polarons in Two-Dimensional Polar Materials

Abstract
We propose a macroscopic theory of optical phonons, Fröhlich polarons, and exciton-polarons in two-dimensional (2D) polar crystalline monolayers. Our theory extends the classical macroscopic formulation of the electron-phonon problem in three-dimensional (3D) polar crystals to the new generation of 2D materials. Similarly to the 3D case, in our approach, the effective electron-phonon Hamiltonian is parametrized solely in terms of macroscopic experimentally accessible quantities – 2D polarizabilities of the monolayer at low and high frequencies. We derive the dispersion of long wave length longitudinal optical (LO) phonons, which can be viewed as a 2D form of the Lyddane-Sachs-Teller relation, and study the formation of 2D Fröhlich polarons by adopting the intermediate coupling approximation. Finally, we apply this approach to excitons in polar 2D crystals and derive an effective potential of the electron-hole interaction dressed by LO phonons. Due to the specific dispersion of LO phonons, the polarons and exciton-polarons in 2D materials exhibit unique features not found in their 3D counterparts. As an illustration, the polaron and exciton-polaron binding energies are computed for a representative set of 2D polar crystals, demonstrating the interplay between dimensionality, polarizability of materials, and electron-phonon coupling.
Polarons, quasiparticles arising from the interaction of charge carriers with lattice vibrations, have played a pivotal role in understanding charge transport and optical properties in polar materials [1]. The concept of polaron was initially formulated as an electron interacting with the induced displacement field in dielectric media, the so-called Landau-Pekar model [2, 3, 4] in which polaronic effects result in autolocalization of the electron in a selfconsistent potential well. Alternatively, a weak and intermediate coupling approach implies a field theoretical picture of the electron dressed by a cloud of longitudinal optical (LO) phonons, known as the Fröhlich polaron [5, 6, 7]. Polaronic effects in bulk polar crystals such as alkali halides and oxides are well studied both theoretically and experimentally [8, 9], providing important benchmarks for strong-coupling electron-phonon physics. Recent advances in this direction include accurate diagrammatic Monte-Carlo results [10, 11], and first-principle calculations [12, 13, 14].
The impact of LO phonons on exciton states leading to the formation of exciton-polarons is widely studied in II-VI and III-V polar semiconductors [15, 16, 17, 18, 19]. Recently, there has been a growing interest in exciton-polaron states in hybrid perovskites [20]. Here, the exciton-polaron concept allows to correctly address the main features of optical spectra both in bulk [21, 22, 23, 24, 25, 26, 27], and layered hybrid perovskites [28, 29, 30, 31, 32, 33, 34].
Despite rapid progress in the fabrication and study of two-dimensional (2D) materials [35], there are a handful of observations of polaronic effects in 2D. This includes hexagonal boron nitride (hBN) [36], MoS2 monolayer [37], SrTiO3 surface [38, 39]. On the theory side, most studies of 2D polarons in different physical situations adopt an ad hoc replacement of bare 3D to 2D Coulomb potential in the electron-phonon Hamiltonian [40, 41, 11, 42, 43]. Recently, a first principle extension of Landau-Pekar polaron theory, which accounts for the proper 2D screening, has been developed [44]. Very recently, the exciton-polaron problem for a single layer GaN was addressed within many-body theory [45, 46] by perturbatively correcting the kernel in the Bethe-Salpeter equation [47].
In 2D insulators, a reduced and highly nonlocal dielectric screening [48, 49, 50] results in a very unusual dispersion of LO phonons [51, 52]. Apparently, the correct treatment of the specific 2D screening is crucial for all phenomena related to LO phonons and electron-phonon interaction in polar 2D systems. Some aspects of this problem have been studied and understood using a thin dielectric slab model combined with ab initio calculations [51, 52, 53]. However, it comes as a surprise that a genuine 2D theory of long wave length LO phonons and electron-phonon interaction in atomically thin polar insulators is still lacking. The present work aims to fill this gap. We propose a macroscopic approach to optical phonon modes in 2D polar materials, which generalizes the classical textbook theory of 3D polar crystals [54]. This provides a unified theoretical framework for describing 2D polarons and exciton-polarons (see Fig. 1 (a)). We derive the electron-phonon Hamiltonian and adopt the intermediate coupling scheme of Lee, Low, and Pines (LLP) [6] to obtain the polaron binding energy and effective mass. We then analyze the exciton problem, where, similarly to the bulk situation [18], we obtain an effective electron-hole interaction potential that includes polaronic corrections. We find that polaron effects substantially modify the exciton binding energy and the shape of wave function, which has implications for interpreting the optical spectra of 2D polar semiconductors.
Figure 1: (a) Schematic of a 2D polar crystal layer hosting an electron-hole pair dressed by LO phonons, leading to the formation of exciton-polaron. (b) Dispersion of LO and TO phonon modes for a hBN monolayer. Here the dielectric constant of surrounding media is set $\varepsilon=1$. The remaining parameters are presented in Table 4.
Phonon dispersion. We consider 2D optical phonon modes in a polar monolayer located at the plane $z=0$. Let us define the mass-weighted displacement field $\bm{\xi}=\sqrt{nM}\mathbf{S}$, where $\mathbf{S}=\mathbf{U}_{+}-\mathbf{U}_{-}$ is the relative displacement of positively and negatively charged ions, $n=N/\Sigma$ is the sheet density of unit cells, and $M=m_{+}m_{-}/(m_{+}+m_{-})$ is the reduced ionic mass. The Lagrangian density of the system reads
$\mathcal{L}=\left[\frac{1}{2}\dot{\bm{\xi}}^{2}-W^{\rm 2D}\right]\delta(z)+\varepsilon_{0}\varepsilon\frac{\mathbf{E}^{2}(\mathbf{r},z)}{2},$ (1)
where $W^{\rm 2D}$ is the 2D potential energy density of the ionic system, $\varepsilon$ is the dielectric constant of the environment, $\varepsilon_{0}$ is the vacuum permittivity, $\mathbf{E}=-\nabla\varphi$ is the electric field, and the last term describes the contribution of electromagnetic field in the bulk environment. Similarly to the 3D case [54], the potential energy within the 2D layer has a form fixed by the symmetry  [^1] For simplicity we assume a 2D isotropy of the layer, but this assumption can be readily relaxed.:
$W^{\rm 2D}=\frac{\omega_{\rm t}^{2}}{2}\bm{\xi}^{2}-\gamma\bm{\xi}\cdot\mathbf{E}(\mathbf{r},z=0)-\frac{\alpha_{\infty}}{2}\mathbf{E}^{2}(\mathbf{r},z=0),$ (2)
where the coefficients are uniquely determined from the high and low frequency asymptotic values of the macroscopic 2D polarization $\mathbf{P}=\partial W^{\rm 2D}/\partial\mathbf{E}$. Specifically, here $\omega_{\rm t}$ is the transverse optical (TO) phonon frequency, $\alpha_{\infty}$ is the high-frequency 2D polarizability, and $\gamma=\omega_{\rm t}\sqrt{\alpha_{0}-\alpha_{\infty}}$ where $\alpha_{0}$ is the static polarizability, see Supplemental Material (SM) for details. Apparently, the difference $\alpha_{0}-\alpha_{\infty}$ characterizes the ionic contribution to the polarizability of the monolayer.
The Euler-Lagrange equations, generated by Eq. 1, consist of the equation of motion (EOM) for the 2D displacement $\bm{\xi}(\mathbf{r})$ and the 3D Poisson equation for the electric potential $\varphi(\mathbf{r},z)$. By considering plane-wave solutions, $\bm{\xi}(\mathbf{r})\sim e^{i\mathbf{k}\cdot\mathbf{r}}$, $\varphi(\mathbf{r},z)\sim e^{i(\mathbf{k}\cdot\mathbf{r}+k_{z}z)}$, and defining the in-plane electric potential in the $\mathbf{k}$-space as $\varphi_{\mathbf{k}}^{\rm 2D}=\int\varphi_{\mathbf{k},k_{z}}{\rm d}k_{z}/(2\pi)$, we find
$\varphi_{\mathbf{k}}^{\rm 2D}=-\frac{i\gamma}{4\pi\varepsilon_{0}k}\cdot\frac{\mathbf{k}\cdot\bm{\xi}_{\mathbf{k}}}{\varepsilon+\alpha_{\infty}k/(2\varepsilon_{0})}.$ (3)
Substitution of this expression into EOM for the displacement, yields the dispersion relation for LO phonons,
$\omega_{{\rm l},k}^{2}=\omega_{\rm t}^{2}{\frac{\varepsilon+r_{0}k}{\varepsilon+r_{\infty}k}},$ (4)
where we introduced the 2D screening lengths as $r_{\infty}=\alpha_{\infty}/(2\varepsilon_{0})$, $r_{0}=\alpha_{0}/(2\varepsilon_{0})$. Equation (4) perfectly agrees the dispersion obtained from heuristic arguments supported by direct first principal calculations in Ref. [52] and recently observed experimentally [56]. Remarkably, in our approach, as expected physically in the long wave length limit, the relation between frequencies of TO an LO phonons involves only experimentally accessible macroscopic polarizabilities. The LO-TO degeneracy at $k=0$ as well as a nontrivial dispersion of LO mode are the manifestations of the dielectric screening in 2D. It is worth noting that in many instances, the quantity $\varepsilon+\alpha k/2\varepsilon_{0}$ enters as an effective 2D dielectric constant, and therefore Eq. (4) can be viewed as a 2D form of the Lyddane-Sachs-Teller relation. We visualize the LO phonon dispersion for hBN monolayer, as depicted in Fig. 1 (b) 222 For the sake of simplicity in all explicit calculations we assume freestanding atomic monolayers, i.e. $\varepsilon=1$. The remaining parameters are summarized in Table 4..
Table 1: Material parameters taken from experimental and first-principal calculations data for a set of 2D polar materials: TO phonon energy $\hbar\omega_{\rm t}$, screening lengths $r_{0}$, $r_{\infty}$444the screening lengths are extracted from Ref. [44] as $r_{0[\infty]}=\epsilon_{0[\infty]}d/2$, where ϵ0[]subscriptitalic-$\epsilon_{0[\infty]}$ is the static [high-frequency] dielectric constant, $d$ is the monolayer thickness., electron and hole effective masses $m_{e}$, $m_{h}$, the bandgap $E_{g}$ (asterisk marks the direct-gap materials). Calculated renormalized masses $m_{e}^{*}=m_{e}(1+\alpha_{e})$, $m_{h}^{*}=m_{h}(1+\alpha_{h})$, polaron binding energies $E^{\rm b}_{e}$, $E^{\rm b}_{h}$, bare exciton binding energy $E_{\rm X}^{0}$, and exciton-polaron binding energy $E_{\rm X}^{\rm b}$. The values of $r_{0}$, $r_{\infty}$, $m_{e}$, $m_{h}$ are taken from Ref. [44].
Material $\hbar\omega_{\rm t}$ (meV) $r_{0}$ (nm) $r_{\infty}$ (nm) $m_{e}/m_{e}^{*}$ $m_{h}/m_{h}^{*}$ $E^{\rm b}_{e}$ (meV) $E^{\rm b}_{h}$ (meV) $E_{\rm X}^{0}/E_{\rm X}^{\rm b}$ (meV) $E_{g}$ (eV)
h-BN $172.3$[52] $1.076$ [52] $0.780$ [52] $0.83$$/1.00$ $0.65$$/0.77$ $129.3$ $116.3$ $1617/1869$ $4.69^{*}$[44]
HfS2 $17.81$[58] $13.98$ $2.974$ $0.24$$/0.54$ $0.44$$/1.07$ $123.2$ $155.1$ $197.8/444.6$ $1.31$ [58]
HfSe2 $11.10$[59] $21.75$ $4.246$ $0.18$$/0.43$ $0.23$$/0.57$ $89.56$ $98.51$ $127.5/314.9$ $0.64$ [59]
GaN $73.30$[60] $1.109$ $0.756$ $0.24$$/0.29$ $1.35$$/2.02$ $55.16$ $132.0$ $1304/1591$ $6.32^{*}$[60]
AlN $74.15$[61] $0.763$ $0.466$ $0.51$$/0.71$ $1.49$$/2.53$ $105.4$ $186.5$ $2064/2891$ $2.81$ [62]
ZrS2 $29.82$[63] $10.62$ $2.826$ $0.31$$/0.58$ $0.26$$/0.48$ $142.9$ $133.7$ $238.5/428.6$ $1.25$ [64]
Figure 2: Screening length ratio dependence of (a) $I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})$, and (b) $I_{m}(\sigma_{0},\sigma_{e,{\rm t}})$ characterizing the polaron binding energy and mass renormalization, respectively. The polar effects become more pronounced with the increase of parameter $\sigma_{0}$, resulting in the enhancement of both $I_{\rm b}$, and $I_{m}$. The reduction of $\sigma_{e,{\rm t}}$ corresponds to increase of phonon energy and [or] free electron effective mass, both resulting in the increase of polaron binding and mass renormalization. Asterisks in panel (a) illustrate the representative polar atomic monolayers.
Fröhlich Hamiltonian and polaron problem. The electric potential of Eq. (3), generated by the ionic displacement, acts on the electrons, generating electron-phonon coupling. Formally, the electron-phonon Hamiltonian is derived by inserting the functional $\varphi[\bm{\xi}_{\mathbf{k}}]$ into the Lagrangian (1) and canonically quantizing the displacement field, see SM. We stress that the last term in Eq. (1) is essential to restore the correct Hamiltonian of LO phonons. The final electron-phonon Hamiltonian reads
$\hat{H}_{e-{\rm ph}}=\frac{\hat{\mathbf{p}}_{e}^{2}}{2m_{e}}+\sum_{\mathbf{k}}\hbar\omega_{{\rm l},k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}}\\ +\sum_{\mathbf{k}}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}^{\dagger}_{\mathbf{k}}\right),$ (5)
where $m_{e}$, $\hat{\mathbf{p}}_{e}$, $\mathbf{r}_{e}$ are the electron mass, momentum, and position, respectively, and $\hat{a}_{\mathbf{k}}$ annihilates a LO phonon with momentum $\mathbf{k}$. The electron-phonon interaction is given by $-e\varphi^{\rm 2D}(\mathbf{r}_{e})$, leading to the coupling amplitude,
$V_{k}=ie\omega_{\rm t}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}\Sigma\omega_{{\rm l},k}}}\frac{1}{\varepsilon+r_{\infty}k},$ (6)
which depends on the LO phonon dispersion and contains a typical 2D screening factor.
The Hamiltonian of Eq. (5) contains three length scales: $r_{0}$, $r_{\infty}$, and $r_{i,{\rm t}}=\sqrt{\hbar/(2\omega_{\rm t}m_{i})}$, $i=e\,[h]$ for electron [hole]. The two ratios, $\sigma_{0}=r_{0}/r_{\infty}$ and $\sigma_{i,{\rm t}}=\varepsilon r_{i,{\rm t}}/r_{\infty}$, define two independent parameters, characterizing the 2D Fröhlich polaron problem, which should be contrasted to the 3D problem determined by a single dimensionless coupling constant.
We analyze the properties of 2D polaron by adopting the LLP intermediate coupling approach [6], which is expected to be accurate in most realistic situations. Firstly, we note the conservation of total momentum, $[\hat{\mathbf{p}}_{e}+\sum_{\mathbf{k}}\hbar\mathbf{k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}},\hat{H}_{e-{\rm ph}}]=0$, and eliminate the electron coordinate by performing a unitary transformation with the operator $\hat{S}=\exp\left(i\left(\mathbf{p}/h-\sum_{\mathbf{k}}\mathbf{k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}}\right)\cdot\mathbf{r}_{e}\right)$. Next, the phonon field is shifted with $\hat{U}=\exp\left(\sum_{\mathbf{k}}\left(f(\mathbf{k})\hat{a}^{\dagger}_{\mathbf{k}}-f^{*}(\mathbf{k})\hat{a}_{\mathbf{k}}\right)\right)$, acting as $\hat{U}^{-1}\hat{a}_{\mathbf{k}}\hat{U}=\hat{a}_{\mathbf{k}}+f(\mathbf{k})$, where $f(\mathbf{k})$ is a $C$-number serving as a variational parameter. The variational anzats reads $|\psi\rangle=\hat{U}|\psi_{0}\rangle$, assuming $\hat{a}_{\mathbf{k}}|\psi_{0}\rangle=0$. The variational minimization leads to the following expression for polaron energy (see SM for the details):
$E_{e}^{\rm pol}\approx-E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})+\frac{p^{2}}{2m_{e}}\frac{1}{1+\alpha_{e}}.$ (7)
Here we introduce the notations $E_{0}=e^{2}/(4\pi\varepsilon_{0}r_{\infty})$, $\alpha_{e}=E_{0}I_{m}(\sigma_{0},\sigma_{e,{\rm t}})/(\hbar\omega_{\rm t})$, and dimensionless quantities
$I_{\rm b}$ $=\int\limits_{0}^{\infty}\frac{(\sigma_{0}-1)x}{2(1+\sigma_{0}x)(1+x)\left[1+\sigma_{e,{\rm t}}^{2}x^{2}\sqrt{\frac{1+x}{1+\sigma_{0}x}}\right]}\,{\rm d}x,$ (8)
$I_{m}$ $=\int\limits_{0}^{\infty}\frac{\sigma_{e,{\rm t}}^{2}(\sigma_{0}-1)x^{3}}{(1+\sigma_{0}x)^{2}\left(1+\sigma_{e,{\rm t}}^{2}x^{2}\sqrt{\frac{1+x}{1+\sigma_{0}x}}\right)^{3}}\,{\rm d}x.$ (9)
The first term in Eq. (7) corresponds to the polaron binding energy $E^{\rm b}_{e}=E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})$, and the second term is the polaron kinetic energy with rescaled effective mass $m_{e}^{*}=m_{e}(1+\alpha_{e})$. Fig. 2 (a), (b) illustrate the factors $I_{\rm b}(\sigma_{0},\sigma_{\rm t})$, $I_{m}(\sigma_{0},\sigma_{\rm t})$ characterizing the polaron binding energy and mass renormalization, respectively. The parameter $\sigma_{0}$ characterizes the strength of polar effects. Therefore, the increase of $\sigma_{0}$ enhances both $I_{\rm b}$ and $I_{m}$. The parameter $\sigma_{e,{\rm t}}\propto 1/\sqrt{\omega_{\rm t}m_{e}}$ is the inverse energy scale of phonon energy and effective mass. Hence, the reduction of $\sigma_{e,{\rm t}}$ enhances both $I_{\rm b}$ and $I_{m}$, in line with the case of bulk polarons, where the polaron binding energy scales as $E_{e}^{\rm pol}\propto\sqrt{\omega_{\rm LO}m_{e}}$ [54]. We recall that the actual values of polaron binding energy and mass renormalization include the material-specific factors $E_{0}$ and $\hbar\omega_{\rm t}$, as it follows from Eq. (7). The respective data for several representative materials is shown in Table 4. The obtained values of polaron binding energy essentially differ from that reported in Ref. [44], where the ab-initio generalization of Landau-Pekar model is developed. We attribute this discrepancy to (i) the strong electron-phonon coupling limit, corresponding to Landau-Pekar model [54, 65], and (ii) the trivial limit of 2D LO phonon dispersion, considered in Refs. [53, 44], which strongly affects the electron–phonon interaction matrix element.
Notably, the resulting expressions for the polaron binding energy and mass renormalization are not of analytical form, unlike the case of bulk polarons [6]. However, substituting $\sqrt{(1+x)/(1+\sigma_{0}x)}\approx 1/\sqrt{\sigma_{0}}$, we get
$I_{\rm b}\approx\frac{\sigma_{0}-1}{2(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})}\left(\frac{\pi(\sigma_{0}+1)\tilde{\sigma}_{e,{\rm t}}}{2\left(\tilde{\sigma}_{e,{\rm t}}^{2}+1\right)}\right.\\ \left.+\frac{\left(\tilde{\sigma}_{e,{\rm t}}^{2}-\sigma_{0}\right)\log(\tilde{\sigma_{e,{\rm t}}})}{\tilde{\sigma}_{e,{\rm t}}^{2}+1}-\frac{\sigma_{0}\log(\sigma_{0})}{\sigma_{0}-1}\right),$ (10)
[…]
Figure 3: Effective potential of electron-hole interaction $V_{\rm eff}$ (solid curve) and its approximate analytical expression $V_{\rm eff}^{\rm an}$ (dashed curve) for (a) h-BN, and (b) HfSe2 monolayer. The dotted curves illustrate the KR potentials $V_{\rm KR}(r,r_{\infty})$ and $V_{\rm KR}(r,r_{0})$.
Figure 4: Spatial profile of exciton wave function for (a) h-BN, and (b) HfSe2 monolayer. Here solid orange curves illustrate the exciton-polaron wave function, and black dotted curves correspond to bare exciton state. (c) Map of 2D polar materials illustrating the relation of single polaron and exciton-polaron binding energies.
Exciton-polaron. When an exciton forms in a 2D polar crystal, both electron and hole couple to LO phonons. The resulting composite quasiparticle can be viewed as an exciton dressed by a phonon cloud, i.e., an exciton-polaron. The Hamiltonian of an electron-hole pair interacting with LO phonons is
$\hat{H}_{\rm X}$ $=\frac{\hat{\mathbf{p}}_{e}^{2}}{2m_{e}}+\frac{\hat{\mathbf{p}}_{h}^{2}}{2m_{h}}+V_{\rm KR}(|\mathbf{r}_{e}-\mathbf{r}_{h}|,r_{\infty})+\sum_{\mathbf{k}}\hbar\omega_{{\rm l},k}\hat{a}^{\dagger}_{\mathbf{k}}\hat{a}_{\mathbf{k}}$
$+\sum_{\mathbf{k}}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}^{\dagger}_{\mathbf{k}}\right)$
$-\sum_{\mathbf{k}}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{h}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{h}}\hat{a}^{\dagger}_{\mathbf{k}}\right),$ (11)
where $V_{\rm KR}$ is the bare 2D Coulomb (Keldysh-Rytova) potential, reading as [48, 49, 50]
$V_{\rm KR}(r,r_{\infty})=-\frac{e^{2}}{4\pi\varepsilon_{0}}\frac{\pi}{2r_{\infty}}\left[H_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)\right],$ (12)
with $H_{0}$, $Y_{0}$ denoting the Struve function, and the Bessel function of the second kind. We transform to center of center-of-mass $\mathbf{R}=(m_{e}\mathbf{r}_{e}+m_{h}\mathbf{r}_{h})/M$, and relative coordinates $\mathbf{r}=\mathbf{r}_{e}-\mathbf{r}_{h}$, where $M=m_{e}+m_{h}$ is the exciton total mass. We then adopt a variational approach developed for bulk exciton-polarons [18]. Particularly, we apply a unitary transformation $\hat{S}_{\rm X}=\exp\left(i\left(\mathbf{Q}-\sum_{k}\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\right)\cdot\mathbf{R}\right)$ which eliminates the center-of-mass coordinate. Here $\hbar\mathbf{Q}$ denotes the total momentum of exciton. Similar to the case of single polaron, we introduce the mean value of displacement field $F_{\mathbf{k}}(\mathbf{r})$, and apply a unitary transformation with operator $\hat{U}_{\rm X}=\exp\left(\sum_{k}\left(\hat{a}_{\mathbf{k}}F^{*}_{\mathbf{k}}(\mathbf{r})-\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{k}}(\mathbf{r})\right)\right)$. Assuming again that the resulting wave function corresponds to a phonon ground-state $\hat{a}_{\mathbf{k}}|\psi_{\rm X}\rangle=0$, and setting the center-of-mas momentum $\hbar\mathbf{Q}=0$, we get the following energy functional (see SM):
$E_{\rm X}$ $=\left\langle\psi_{\rm X}\left|\hat{\mathbf{p}}^{2}/(2\mu)+V_{\rm KR}(r,r_{\infty})+W_{\rm X}(\mathbf{r})\right|\psi_{\rm X}\right\rangle,$ (13)
where $\mu=m_{e}m_{h}/(m_{e}+m_{h})$ is the exciton reduced mass,
$W_{\rm X}(\mathbf{r})=\frac{\hbar^{2}}{2\mu}\sum_{k}\nabla F^{*}_{\mathbf{k}}\cdot\nabla F_{\mathbf{k}}+\sum_{k}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)|F_{\mathbf{k}}|^{2}\\ -\sum_{k}\left(V_{k}\rho_{\mathbf{k}}F_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}F_{\mathbf{k}}^{*}\right),$ (14)
and $\rho_{\mathbf{k}}=\exp\left(im_{h}\mathbf{k}\cdot\mathbf{r}/M\right)-\exp\left(-im_{e}\mathbf{k}\cdot\mathbf{r}/M\right)$. Unlike the case of single polarons, here the presence of the relative coordinate and relative momentum operator $\hat{\mathbf{p}}$ couples the displacement field $F_{\mathbf{k}}(\mathbf{r})$ and the exciton wave function $\psi_{\rm X}$. For the sake of simplicity, we ignore this mixing, and apply the variational minimization $\partial W_{\rm X}/\partial F_{\mathbf{k}}=0$. The result reads as follows (see SM):
$E_{\rm X}^{\rm b}$ $=\left\langle\psi_{\rm X}\left|\hat{\mathbf{p}}^{2}/(2\mu^{*})+V_{\rm eff}(r)\right|\psi_{\rm X}\right\rangle.$ (15)
Here we defined the exciton binding energy as $E_{\rm X}^{\rm b}=E_{\rm X}+E^{\rm b}_{e}+E^{\rm b}_{h}$, and phenomenologically included the mass renormalizations: $\mu^{*}=m_{e}^{*}m_{h}^{*}/(m_{e}^{*}+m_{h}^{*})$. Such renormalization naturally emerges in the full treatment [19]. The effective Coulomb potential can be written as
$V_{\rm eff}^{\rm ap}(r)$ $\approx V_{\rm KR}(r,r_{0})$
$-E_{0}\left[\frac{m_{h}}{\Delta m}h_{\tilde{\sigma}_{h,t}}\left(\frac{\varepsilon r}{r_{\infty}}\right)-\frac{m_{e}}{\Delta m}h_{\tilde{\sigma}_{e,t}}\left(\frac{\varepsilon r}{r_{\infty}}\right)\right],$ (16)
where $\Delta m=m_{h}-m_{e}$, and
$h_{\nu}(x)=\frac{\sigma_{0}-1}{2(1+\nu^{2})(\sigma_{0}^{2}+\nu^{2})}\left(2(\nu^{2}-\sigma_{0})K_{0}\left(x/\nu\right)\right.$
$\left.+(1+\sigma_{0})\alpha\pi\left[I_{0}\left(x/\nu\right)-L_{0}\left(x/\nu\right)\right]\right)$ (17)
is the function expressing the phonon-induced corrections. Here $I_{0}$, $K_{0}$ are the modified Bessel functions of the first and the second kind, respectively, and $L_{0}$ is the modified Struve function. The Eq. (16) was derived in an approximation $\tilde{\sigma}_{e,{\rm t}}^{2},\tilde{\sigma}_{h,{\rm t}}^{2}\ll 1$, and the total expression is given in SM.
The modified Coulomb potential given by Eq. (16) is the 2D counterpart of Pollmann-Büttner potential, characterizing the bulk exciton-polarons [18]. Fig. 3 (a), (b) illustrate the coordinate dependence of effective potential for hBN and HfSe2 monolayers, respectively. Near the origin $V_{\rm eff}\rightarrow V_{\rm KR}(r,r_{\infty})$, whereas for the long-range tail $V_{\rm eff}\rightarrow V_{\rm KR}(r,r_{0})$, analogous to the case of bulk polarons [20]. The polaronic modification of the effective potential is more pronounced for HfSe2 monolayer shown in Fig. 3 (b), which is due to strong polar effects characterized by the condition $\sigma_{0}\gg 1$.
Using the obtained effective potential, we evaluate the exciton-polaron binding energy, defined by Eq. (15). Figures 4 (a), (b) illustrate the exciton wave functions in hBN and HfSe2 monolayers, respectively. The dotted curves correspond to bare exciton states, discarding the polaron effects. The drastic reshaping of the wave function in the HfSe2 monolayer is again the manifestation of strong polaronic effects. This also translates to the values of exciton binding energy (314.9 meV vs 127.5 meV for bare exciton). Here the binding energy of the bare exciton $E_{\rm X}^{0}$ is calculated with the statically screened potential $V_{\rm KR}(r,r_{0})$. For hBN monolayer the polaron effects are less pronounced, and the exciton-polaron binding energy is 1869.2 meV (1616 meV for bare exciton). Fig. 4 (c) summarizes both the single polaron and exciton-polaron binding energy data for a range of polar atomic monolayers. Here $\Delta E_{\rm X}^{\rm b}=E_{\rm X}^{\rm b}-E_{\rm X}^{0}$ characterizes the modulation of exciton binding energy stemming from polaron effects.
Conclusion. We developed a variational theory of polaron and exciton-polaron formation in 2D polar materials. Our macroscopic model allows to derive an analytical expression for LO phonon dispersion via simple electrostatic considerations, which reproduces the results of first-principal calculations [52]. Variational treatment of electron-phonon coupling problem results in semi-analytical expressions for the binding energy and mass renormalization of Frohlich polaron, which essentially differs from its bulk counterpart [6]. The extension of variational method for the exciton-polaron problem reveals the modulation of pure 2D electron-hole Coulomb interaction due to the coupling of charge carriers to phonons, in line with the case of bulk exciton-polarons [18]. The framework of pure 2D polarons presented here can be readily expanded to account for the finite thickness, characteristic to layered hybrid perovskites [66].
SUPPLEMENTAL MATERIAL for:
Polarons and Exciton-Polarons in Two-Dimensional Polar Materials
V. Shahnazaryan1, A. Kudlis2,1, and I. Tokatly3,4,5
1Abrikosov Center for Theoretical Physics, Dolgoprudnyi, Moscow Region 141701, Russia
2Science Institute, University of Iceland, Dunhagi 3, IS-107, Reykjavik, Iceland
3Departamento de Polímeros y Materiales Avanzados: Física,
Química y Tecnología, Universidad del País Vasco,
Avenida Tolosa 72, E-20018 San Sebastián, Spain
4Donostia International Physics Center (DIPC), E-20018 Donostia-San Sebastián, Spain
5IKERBASQUE, Basque Foundation for Science, Plaza Euskadi 5, 48009 Bilbao, Spain
I 2D LO phonon dispersion
I.1 Lagrangian density
The Lagrangian density of the system is given by:
$\mathcal{L}=\left(\frac{1}{2}\dot{\bm{\xi}}^{2}-W^{\rm 2D}\right)\delta(z)+\varepsilon_{0}\varepsilon\dfrac{\mathbf{E}^{2}}{2}.$ (S1)
Here the two terms in brackets correspond to surface density of kinetic and potential energies, respectively. The last term is the electromagnetic energy density of the surrounding media. The potential energy density accounts for the energy of ionic displacements, and their interaction with the electric field $\mathbf{E}$:
$W^{\rm 2D}$ $=\frac{1}{2}\omega_{\rm t}^{2}\bm{\xi}^{2}-\gamma\bm{\xi}\cdot\mathbf{E}(\mathbf{r},z=0)-\frac{\alpha_{\infty}}{2}\mathbf{E}^{2}(\mathbf{r},z=0).$ (S2)
The 2D polarization field $\mathbf{P}^{\rm 2D}$ then reads:
$\mathbf{P}^{\rm 2D}$ $=-\dfrac{\partial W^{\rm 2D}}{\partial\mathbf{E}}=\gamma\bm{\xi}+\alpha_{\infty}\mathbf{E}.$ (S3)
Taking the derivative of the potential energy with respect to the displacement field, we obtain:
$\frac{\partial W^{\rm 2D}}{\partial\bm{\xi}}$ $=\omega_{\rm t}^{2}\bm{\xi}-\gamma\mathbf{E}.$ (S4)
In the static case ($\omega=0$), the system is in equilibrium, so the derivative of the potential energy with respect to $\bm{\xi}$ must vanish:
$\frac{\partial W^{\rm 2D}}{\partial\bm{\xi}}$ $=0\implies\omega_{\rm t}^{2}\bm{\xi}_{0}-\gamma\mathbf{E}=0.$ (S5)
Solving with respect to the equilibrium displacement $\bm{\xi}_{0}$, we find:
$\bm{\xi}_{0}$ $=\frac{\gamma}{\omega_{\rm t}^{2}}\mathbf{E},$ (S6)
which gives the polarization $\mathbf{P}_{0}^{\rm 2D}$ in equilibrium
$\mathbf{P}_{0}^{\rm 2D}$ $=\gamma\bm{\xi}_{0}+\alpha_{\infty}\mathbf{E}=\left(\alpha_{\infty}+\frac{\gamma^{2}}{\omega_{\rm t}^{2}}\right)\mathbf{E}=\alpha_{0}\mathbf{E},$ (S7)
where we have defined the static dielectric polarizability $\alpha_{0}$ as:
$\alpha_{0}$ $=\alpha_{\infty}+\frac{\gamma^{2}}{\omega_{\rm t}^{2}}.$ (S8)
Expressing the coupling constant $\gamma$ in terms of $\omega_{t}$, $\alpha_{0}$, and $\alpha_{\infty}$, we have:
$\gamma$ $=\omega_{\rm t}\sqrt{\alpha_{0}-\alpha_{\infty}}.$ (S9)
I.2 Equations of Motion (EOM)
The equation of motion for the displacement field $\bm{\xi}$ is derived from the Lagrangian:
$\ddot{\bm{\xi}}$ $=-\omega_{\rm t}^{2}\bm{\xi}+\gamma\mathbf{E}(\mathbf{r},z=0).$ (S10)
Here, the electric field $\mathbf{E}$ depends on the two-dimensional coordinate vector $\mathbf{r}$, and the condition $z=0$ is due to the two-dimensional nature of the system. Gauss’s law relates the divergence of the electric field to the charge density $\rho$:
$\nabla\cdot\mathbf{E}({\mathbf{r}},z)$ $=\frac{\rho}{\varepsilon\varepsilon_{0}}\equiv-\frac{1}{\varepsilon\varepsilon_{0}}\nabla\cdot\mathbf{P}^{\rm 2D}\delta(z).$ (S11)
I.3 Dispersion relation
Expressing the electric field in terms of the scalar potential $\varphi$
$\mathbf{E}$ $=-\nabla\varphi,$ (S12)
and substituting into the equations of motion, we get:
$\ddot{\bm{\xi}}\delta(z)$ $=\left(-\omega_{\rm t}^{2}\bm{\xi}-\gamma\nabla\varphi|_{z=0}\right)\delta(z),$ (S13)
$\nabla^{2}\varphi(\bm{r},z)$ $=\frac{1}{\varepsilon\varepsilon_{0}}\nabla\cdot\left(\gamma\bm{\xi}-\alpha_{\infty}\nabla\varphi|_{z=0}\right)\delta(z).$ (S14)
We seek for plane-wave solutions for the potential and displacement fields:
$\varphi(\mathbf{r},z)$ $\sim e^{i(\mathbf{q}\cdot\mathbf{r}+k_{z}z)},\quad\xi(\mathbf{r})\sim e^{i\mathbf{q}\cdot\mathbf{r}}$ (S15)
where $\mathbf{q}$, $k_{z}$ are the in-plane and normal-to-plane components of the wavevector. Applying the Fourier transform, we find:
$-(q^{2}+k_{z}^{2})\varphi_{\mathbf{q},k_{z}}$ $=\frac{1}{\varepsilon\varepsilon_{0}}i\mathbf{q}\left(\gamma\bm{\xi}_{\mathbf{q}}/(2\pi)-i\alpha_{\infty}\mathbf{q}\varphi_{\mathbf{q}}^{\rm 2D}\right),$ (S16)
where
$\varphi_{\mathbf{q}}^{\rm 2D}=\int\varphi_{\mathbf{q},k_{z}}\frac{{\rm d}k_{z}}{2\pi},$ (S17)
$\bm{\xi}_{\mathbf{q}}=\frac{1}{(2\pi)^{2}}\int e^{-i\mathbf{q}\mathbf{r}}\bm{\xi}(\mathbf{r}){\rm d^{2}}\mathbf{r}.$ (S18)
Solving with respect to the potential $\varphi_{\mathbf{q},k_{z}}$
$\varphi_{\mathbf{q},k_{z}}$ $=-\frac{1}{\varepsilon\varepsilon_{0}}\frac{1}{q^{2}+k_{z}^{2}}\left(i\gamma\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}/(2\pi)+\alpha_{\infty}q^{2}\varphi_{\mathbf{q}}^{\rm 2D}\right),$ (S19)
and plugging in, we reach at:
$\varphi_{\mathbf{q}}^{\rm 2D}$ $=-\frac{1}{2q\varepsilon\varepsilon_{0}}\left(i\gamma\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}/(2\pi)+\alpha_{\infty}q^{2}\varphi_{\mathbf{q}}^{\rm 2D}\right),$ (S20)
finally yielding in
$\varphi_{\mathbf{q}}^{\rm 2D}$ $=-\frac{1}{4\pi\varepsilon\varepsilon_{0}}\frac{\gamma}{1+\alpha_{\infty}q/(2\varepsilon\varepsilon_{0})}i\frac{\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}}{q}=-\frac{1}{4\pi\varepsilon\varepsilon_{0}}\frac{\gamma}{1+\alpha_{\infty}q/(2\varepsilon\varepsilon_{0})}i{\xi}_{q}^{\rm l},$ (S21)
where we define the longitudinal component of the displacement field:
$\xi_{q}^{\rm l}$ $=\frac{\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}}{q}.$ (S22)
We then substitute back into the equation of motion (S13), and perform Fourier transform, reaching at
$\frac{1}{2\pi}\omega_{\rm l}^{2}\xi_{q,\omega}^{\rm l}$ $=\frac{1}{2\pi}\omega_{\rm t}^{2}\xi_{q,\omega}^{\rm l}+i\gamma q\varphi_{\mathbf{q}}^{\rm 2D}.$ (S23)
We thus find the dispersion relation for the longitudinal optical phonon modes:
$\omega_{{\rm l},q}^{2}$ $=\omega_{\rm t}^{2}+\frac{1}{2\varepsilon\varepsilon_{0}}\frac{\gamma^{2}q}{1+\alpha_{\infty}q/(2\varepsilon\varepsilon_{0})}.$ (S24)
Substituting the expression for $\gamma$, we get:
$\omega_{{\rm l},q}^{2}$ $=\omega_{\rm t}^{2}+\frac{1}{2\varepsilon\varepsilon_{0}}\frac{\omega_{\rm t}^{2}(\alpha_{0}-\alpha_{\infty})q}{1+\alpha_{\infty}q/(2\varepsilon\varepsilon_{0})}.$ (S25)
Therefore, the frequency of the longitudinal optical phonons is:
$\omega_{{\rm l},q}^{2}$ $=\omega_{\rm t}^{2}\frac{1+\alpha_{0}q/(2\varepsilon\varepsilon_{0})}{1+\alpha_{\infty}q/(2\varepsilon\varepsilon_{0})}=\omega_{\rm t}^{2}\frac{\varepsilon+r_{0}q}{\varepsilon+r_{\infty}q}.$ (S26)
II LO phonon Hamiltonian
We introduce generalized momentum as
$\bm{\eta}=\frac{\mathcal{L}}{\partial\dot{\bm{\xi}}}=\dot{\bm{\xi}}\delta(z).$ (S27)
The Hamiltonian density then reads
$\mathcal{H}=\dot{\bm{\xi}}\bm{\eta}-\mathcal{L}=\left(\frac{\dot{\bm{\xi}}^{2}}{2}+\frac{1}{2}\omega_{\rm t}^{2}\bm{\xi}^{2}-\gamma\bm{\xi}\cdot\mathbf{E}(\mathbf{r},z=0)-\frac{\alpha_{\infty}}{2}\mathbf{E}^{2}(\mathbf{r},z=0)\right)\delta(z)-\varepsilon_{0}\varepsilon\frac{\mathbf{E}^{2}}{2},$ (S28)
resulting in the classical Hamiltonian
$H=\int\mathcal{H}{\rm d}^{3}r=\int\left[\left(\frac{\dot{\bm{\xi}}^{2}}{2}+\frac{1}{2}\omega_{\rm t}^{2}\bm{\xi}^{2}+\gamma\bm{\xi}\cdot\nabla\varphi(\mathbf{r},z=0)-\frac{\alpha_{\infty}}{2}\nabla(\varphi(\mathbf{r},z=0))^{2}\right)\delta(z)-\varepsilon_{0}\varepsilon\frac{(\nabla\varphi)^{2}}{2}\right]{\rm d^{2}}r{\rm d}z.$ (S29)
We evaluate the electric field as
$\nabla\varphi(\mathbf{r},z=0)$ $=\nabla\left[\int e^{i\mathbf{q}\cdot\mathbf{r}}\varphi_{\mathbf{q},k_{z}}{\rm d^{2}}q{\rm d}k_{z}\right]=2\pi i\int\mathbf{q}e^{i\mathbf{q}\cdot\mathbf{r}}\frac{-i}{4\pi\varepsilon_{0}}\frac{\gamma}{\varepsilon+r_{\infty}q}\frac{\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}}{q}{\rm d^{2}}q=\frac{\gamma}{2\varepsilon_{0}}\int\frac{q}{\varepsilon+r_{\infty}q}e^{i\mathbf{q}\cdot\mathbf{r}}\bm{\xi}_{\mathbf{q}}^{||}{\rm d^{2}}q.$ (S30)
Here in the last relation we introduce a definition $\bm{\xi}_{\mathbf{q}}^{||}=\xi_{\mathbf{q}}^{\rm l}\mathbf{q}/q=(\bm{\xi}_{\mathbf{q}}\cdot\mathbf{q})\mathbf{q}/q^{2}$. Similarly,
$\nabla\varphi(\mathbf{r},z)$ $=\nabla\left[\int e^{i(\mathbf{q}\cdot\mathbf{r}+k_{z}z)}\varphi_{\mathbf{q},k_{z}}{\rm d^{2}}q{\rm d}k_{z}\right]=i\int(\mathbf{q}+\hat{\mathbf{k}}_{z}k_{z})e^{i(\mathbf{q}\cdot\mathbf{r}+k_{z}z)}\varphi_{\mathbf{q},k_{z}}{\rm d}k_{z}{\rm d^{2}}q$
$=i\int(\mathbf{q}+\hat{\mathbf{k}}_{z}k_{z})e^{i(\mathbf{q}\cdot\mathbf{r}+k_{z}z)}\frac{(-1)}{\varepsilon\varepsilon_{0}}\frac{1}{q^{2}+k_{z}^{2}}\left(i\gamma\mathbf{q}\cdot\frac{\bm{\xi}_{\mathbf{q}}}{2\pi}+\alpha_{\infty}q^{2}\varphi_{\mathbf{q}}^{\rm 2D}\right){\rm d}k_{z}{\rm d^{2}}q$
$=\frac{\gamma}{2\pi\varepsilon_{0}}\int e^{i(\mathbf{q}\cdot\mathbf{r}+k_{z}z)}\frac{(\mathbf{q}+\hat{\mathbf{k}}_{z}k_{z})}{q^{2}+k_{z}^{2}}\frac{1}{\varepsilon+r_{\infty}q}\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}{\rm d}k_{z}{\rm d^{2}}q.$ (S31)
We recall that $\bm{\xi}(\mathbf{r})=\sqrt{MN/\Sigma}\mathbf{S}(\mathbf{r})$, and introduce a plane-wave expansion
$\mathbf{S}(\mathbf{r})=\frac{1}{\sqrt{N}}\sum_{k}e^{i\mathbf{k}\cdot\mathbf{r}}\mathbf{X}_{\mathbf{k}},$ (S32)
and evaluate the terms one by one. Namely,
$\frac{\omega_{\rm t}^{2}}{2}\int\bm{\xi}^{2}(\mathbf{r})\delta(z){\rm d^{2}}r{\rm d}z$ $=\frac{\omega_{\rm t}^{2}}{2}\frac{MN}{\Sigma N}\int\sum_{k}\sum_{k^{\prime}}e^{i(\mathbf{k}+\mathbf{k}^{\prime})\cdot\mathbf{r}}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{\mathbf{k}^{\prime}}{\rm d^{2}}r=\frac{M\omega_{\rm t}^{2}}{2}\sum_{k}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{-\mathbf{k}},$ (S33)
$\frac{1}{2}\int\dot{\bm{\xi}}^{2}(\mathbf{r})\delta(z){\rm d^{2}}r{\rm d}z$ $=\frac{1}{2}\frac{MN}{\Sigma N}\int\sum_{k}\sum_{k^{\prime}}e^{i(\mathbf{k}+\mathbf{k}^{\prime})\cdot\mathbf{r}}\dot{\mathbf{X}}_{\mathbf{k}}\cdot\dot{\mathbf{X}}_{\mathbf{k}^{\prime}}{\rm d^{2}}r=\frac{M}{2}\sum_{k}\dot{\mathbf{X}}_{\mathbf{k}}\cdot\dot{\mathbf{X}}_{-\mathbf{k}}=\frac{1}{2M}\sum_{k}\mathbf{P}_{\mathbf{k}}\cdot\mathbf{P}_{-\mathbf{k}},$ (S34)
where $\mathbf{P}_{\mathbf{k}}=M\dot{\mathbf{X}}_{\mathbf{k}}$. Noting that $\bm{\xi}_{\mathbf{k}}=\mathbf{X}_{\mathbf{k}}\sqrt{\Sigma M}/(2\pi)^{2}$, we get
$\gamma\int\bm{\xi}(\mathbf{r})\cdot\nabla\varphi(\mathbf{r},z=0)\delta(z){\rm d^{2}}r{\rm d}z$ $=\gamma\int\bm{\xi}(\mathbf{r})\cdot\frac{\gamma}{2\varepsilon_{0}}\int\frac{q}{\varepsilon+r_{\infty}q}e^{i\mathbf{q}\cdot\mathbf{r}}\bm{\xi}_{\mathbf{q}}^{||}\delta(z){\rm d^{2}}r{\rm d}z{\rm d^{2}}q$
$=\frac{\gamma^{2}}{2\varepsilon_{0}}\sqrt{\frac{MN}{\Sigma}}\frac{1}{\sqrt{N}}\int\sum_{k}e^{i\mathbf{k}\cdot\mathbf{r}}\frac{q}{\varepsilon+r_{\infty}q}e^{i\mathbf{q}\cdot\mathbf{r}}\frac{\sqrt{\Sigma M}}{(2\pi)^{2}}\mathbf{X}_{\mathbf{q}}^{||}\cdot\mathbf{X}_{\mathbf{k}}{\rm d^{2}}r{\rm d^{2}}q$
$=M\omega_{\rm t}^{2}\sum_{k}\frac{(r_{0}-r_{\infty})k}{\varepsilon+r_{\infty}k}\mathbf{X}_{\mathbf{k}}^{||}\cdot\mathbf{X}_{-\mathbf{k}}^{||}$ (S35)
$\frac{\alpha_{\infty}}{2}\int\left[\nabla\varphi(\mathbf{r},z=0)\right]^{2}\delta(z){\rm d^{2}}r{\rm d}z$ $=\frac{\alpha_{\infty}}{2}\int\frac{\gamma^{2}}{(2\varepsilon_{0})^{2}}\frac{q}{\varepsilon+r_{\infty}q}\frac{k}{\varepsilon+r_{\infty}k}e^{i(\mathbf{q}+\mathbf{k})\cdot\mathbf{r}}\bm{\xi}_{\mathbf{q}}^{||}\cdot\bm{\xi}_{\mathbf{k}}^{||}\delta(z){\rm d^{2}}r{\rm d}z{\rm d^{2}}q{\rm d^{2}}k$
$=\frac{\alpha_{\infty}}{2}\frac{\gamma^{2}}{(2\varepsilon_{0})^{2}}\int\frac{q}{\varepsilon+r_{\infty}q}\frac{k}{\varepsilon+r_{\infty}k}\delta(\mathbf{q}+\mathbf{k})(2\pi)^{2}\frac{\Sigma M}{(2\pi)^{4}}\mathbf{X}_{\mathbf{q}}^{||}\cdot\mathbf{X}_{\mathbf{k}}^{||}{\rm d^{2}}q{\rm d^{2}}k$
$=\frac{M}{2}\omega_{\rm t}^{2}\sum_{k}\frac{(r_{0}-r_{\infty})kr_{\infty}k}{(\varepsilon+r_{\infty}k)^{2}}\mathbf{X}_{\mathbf{-k}}^{||}\cdot\mathbf{X}_{\mathbf{k}}^{||}$ (S36)
$\frac{\varepsilon\varepsilon_{0}}{2}\int(\nabla\varphi)^{2}{\rm d}^{2}r{\rm d}z$
$=\frac{\varepsilon\varepsilon_{0}}{2}\frac{1}{(2\pi)^{2}}\frac{\gamma^{2}}{\varepsilon_{0}^{2}}\int e^{i(\mathbf{q}\cdot\mathbf{r}+q_{z}z)}e^{i(\mathbf{k}\cdot\mathbf{r}+k_{z}z)}\frac{\mathbf{q}+\hat{\mathbf{k}}_{z}q_{z}}{q^{2}+q_{z}^{2}}\cdot\frac{\mathbf{k}+\hat{\mathbf{k}}_{z}k_{z}}{k^{2}+k_{z}^{2}}\frac{1}{\varepsilon+r_{\infty}q}\frac{1}{\varepsilon+r_{\infty}k}(\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}})(\mathbf{k}\cdot\bm{\xi}_{\mathbf{k}}){\rm d}q_{z}{\rm d^{2}}q{\rm d}k_{z}{\rm d^{2}}k{\rm d}^{2}r{\rm d}z$
$=\frac{2\pi\varepsilon\gamma^{2}}{2\varepsilon_{0}}\int\delta(\mathbf{q}+\mathbf{k})\delta(k_{z}+q_{z})\frac{\mathbf{q}\cdot\mathbf{k}+q_{z}k_{z}}{(q^{2}+q_{z}^{2})(k^{2}+k_{z}^{2})}\frac{1}{\varepsilon+r_{\infty}q}\frac{1}{\varepsilon+r_{\infty}k}(\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}})(\mathbf{k}\cdot\bm{\xi}_{\mathbf{k}}){\rm d}q_{z}{\rm d^{2}}q{\rm d}k_{z}{\rm d^{2}}k$
$=\frac{2\pi\varepsilon\gamma^{2}}{2\varepsilon_{0}}\int\frac{k^{2}+k_{z}^{2}}{(k^{2}+k_{z}^{2})^{2}}\frac{1}{(\varepsilon+r_{\infty}k)^{2}}k^{2}\bm{\xi}_{\mathbf{-k}}^{||}\cdot\bm{\xi}_{\mathbf{k}}^{||}{\rm d}k_{z}{\rm d^{2}}k$
$=\frac{M}{2}\omega_{\rm t}^{2}\varepsilon\sum_{k}\frac{(r_{0}-r_{\infty})k}{(\varepsilon+r_{\infty}k)^{2}}\mathbf{X}_{\mathbf{-k}}^{||}\cdot\mathbf{X}_{\mathbf{k}}^{||}$ (S37)
Collecting all the terms, we reach at
$H$ $=\frac{1}{2M}\sum_{k}\mathbf{P}_{\mathbf{k}}\cdot\mathbf{P}_{-\mathbf{k}}+\frac{M}{2}\omega_{\rm t}^{2}\sum_{k}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{-\mathbf{k}}+M\omega_{\rm t}^{2}\sum_{k}\mathbf{X}_{\mathbf{k}}^{||}\cdot\mathbf{X}_{-\mathbf{k}}^{||}\left[\frac{(r_{0}-r_{\infty})k}{\varepsilon+r_{\infty}k}-\frac{1}{2}\frac{(r_{0}-r_{\infty})kr_{\infty}k}{(\varepsilon+r_{\infty}k)^{2}}-\frac{\varepsilon}{2}\frac{(r_{0}-r_{\infty})k}{(\varepsilon+r_{\infty}k)^{2}}\right]$
$=\frac{1}{2M}\sum_{k}\mathbf{P}_{\mathbf{k}}\cdot\mathbf{P}_{-\mathbf{k}}+\frac{M}{2}\omega_{\rm t}^{2}\sum_{k}\mathbf{X}_{\mathbf{k}}\cdot\mathbf{X}_{-\mathbf{k}}+M\omega_{\rm t}^{2}\sum_{k}\mathbf{X}_{\mathbf{k}}^{||}\cdot\mathbf{X}_{-\mathbf{k}}^{||}\frac{(r_{0}-r_{\infty})k}{\varepsilon+r_{\infty}k}\left(1-\frac{1}{2}\frac{r_{\infty}k}{\varepsilon+r_{\infty}k}-\frac{\varepsilon}{2}\frac{1}{\varepsilon+r_{\infty}k}\right)$
$=\sum_{k}\left(\frac{1}{2M}\mathbf{P}_{\mathbf{k}}^{\perp}\cdot\mathbf{P}_{-\mathbf{k}}^{\perp}+\frac{M\omega_{{\rm t}}^{2}}{2}\mathbf{X}_{\mathbf{k}}^{\perp}\cdot\mathbf{X}_{-\mathbf{k}}^{\perp}\right)+\sum_{k}\left(\frac{1}{2M}\mathbf{P}_{\mathbf{k}}^{||}\cdot\mathbf{P}_{-\mathbf{k}}^{||}+\frac{M\omega_{{\rm l},k}^{2}}{2}\mathbf{X}_{\mathbf{k}}^{||}\cdot\mathbf{X}_{-\mathbf{k}}^{||}\right)=H_{\rm TO}+H_{\rm LO},$ (S38)
where $\mathbf{\xi}_{\mathbf{k}}^{\perp}=\mathbf{\xi}_{\mathbf{k}}-\mathbf{\xi}_{\mathbf{k}}^{||}$ is the transverse component of the displacement.
III Electron-phonon coupling
Introducing the notations
$\mathbf{X}_{\mathbf{k}}^{||}$ $=\sqrt{\frac{\hbar}{2M\omega_{{\rm l},k}}}\left(a_{\mathbf{k}}+a_{-\mathbf{k}}^{\dagger}\right)\mathbf{e}_{\mathbf{k}},$ (S39)
$\mathbf{P}_{\mathbf{k}}^{||}$ $=\frac{1}{i}\sqrt{\frac{\hbar M\omega_{{\rm l},k}}{2}}\left(a_{\mathbf{k}}-a_{-\mathbf{k}}^{\dagger}\right)\mathbf{e}_{\mathbf{k}},$ (S40)
We can rewrite the classical Hamiltonian in the form
$H_{\rm LO}=\sum_{k}\frac{\hbar\omega_{{\rm l},k}}{2}(a_{\mathbf{k}}a_{\mathbf{k}}^{\dagger}+a_{-\mathbf{k}}^{\dagger}a_{-\mathbf{k}})$ (S41)
The longitudinal displacement then takes a form
$\mathbf{S}^{||}(\mathbf{r})$ $=\frac{1}{\sqrt{N}}\sum_{k}e^{i\mathbf{k}\cdot\mathbf{r}}\mathbf{X}_{\mathbf{k}}^{||}=\frac{1}{\sqrt{N}}\sum_{k}\sqrt{\frac{\hbar}{2M\omega_{{\rm l},k}}}(e^{i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}+e^{i\mathbf{k}\cdot\mathbf{r}}a_{-\mathbf{k}}^{\dagger})\mathbf{e}_{\mathbf{k}}$
$=\frac{1}{\sqrt{N}}\sum_{k}\sqrt{\frac{\hbar}{2M\omega_{{\rm l},k}}}\left(e^{i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}+e^{-i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}^{\dagger}\right)\mathbf{e}_{\mathbf{k}},$ (S42)
where in the last relation we used the fact $\mathbf{e}_{-\mathbf{k}}=\mathbf{e}_{\mathbf{k}}$ for the given (longitudinal) polarization of displacement field. Correspondingly, the mass-weighted displacement is
$\bm{\xi}^{||}(\mathbf{r})$ $=\sqrt{\frac{MN}{\Sigma}}\mathbf{S}^{||}(\mathbf{r})=\sum_{k}\sqrt{\frac{\hbar}{2\Sigma\omega_{{\rm l},k}}}\left(e^{i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}+e^{-i\mathbf{k}\cdot\mathbf{r}}a_{\mathbf{k}}^{\dagger}\right)\mathbf{e}_{\mathbf{k}}.$ (S43)
The interaction of charge carrier with the electric field by LO phonon is $-e\varphi(\mathbf{r}_{e},0)$, where the LO phonon potential is readily evaluated as
$\varphi(\mathbf{r},0)$ $=\int e^{i\mathbf{q}\cdot\mathbf{r}}\varphi_{\mathbf{q},k_{z}}{\rm d}^{2}q{\rm d}k_{z}=\int e^{i\mathbf{q}\cdot\mathbf{r}}\frac{(-1)}{\varepsilon\varepsilon_{0}}\frac{1}{q^{2}+k_{z}^{2}}\left(i\gamma\mathbf{q}\cdot\frac{\bm{\xi}_{\mathbf{q}}}{2\pi}+\alpha_{\infty}q^{2}\varphi_{\mathbf{q}}^{\rm 2D}\right){\rm d}^{2}q{\rm d}k_{z}$
$=-\frac{1}{\varepsilon\varepsilon_{0}}\int e^{i\mathbf{q}\cdot\mathbf{r}}\frac{1}{q^{2}+k_{z}^{2}}\left(i\gamma\mathbf{q}\cdot\frac{\bm{\xi}_{\mathbf{q}}}{2\pi}-\alpha_{\infty}q^{2}\frac{1}{4\pi\varepsilon_{0}}\frac{\gamma}{\varepsilon+r_{\infty}q}i\frac{\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}}{q}\right){\rm d}^{2}q{\rm d}k_{z}$
$=-\frac{i\gamma}{2\varepsilon_{0}}\int e^{i\mathbf{q}\cdot\mathbf{r}}\frac{1}{q}\frac{1}{\varepsilon+r_{\infty}q}\mathbf{q}\cdot\bm{\xi}_{\mathbf{q}}{\rm d}^{2}q.$ (S44)
Further,
$\bm{\xi}_{\mathbf{q}}^{||}=\frac{\sqrt{\Sigma M}}{(2\pi)^{2}}\mathbf{X}_{\mathbf{q}}^{||}=\sqrt{\frac{\hbar\Sigma}{2\omega_{{\rm l},q}}}\frac{1}{(2\pi)^{2}}\left(a_{\mathbf{q}}+a_{-\mathbf{q}}^{\dagger}\right)\mathbf{e}_{\mathbf{q}},$ (S45)
so that
$\varphi(\mathbf{r},0)$ $=-\frac{i\gamma}{2\varepsilon_{0}}\int e^{i\mathbf{q}\cdot\mathbf{r}}\frac{1}{q}\frac{1}{\varepsilon+r_{\infty}q}\mathbf{q}\cdot\sqrt{\frac{\hbar\Sigma}{2\omega_{{\rm l},q}}}\frac{1}{(2\pi)^{2}}\left(a_{\mathbf{q}}+a_{-\mathbf{q}}^{\dagger}\right)\mathbf{e}_{\mathbf{q}}{\rm d}^{2}q$
$=-i\omega_{\rm t}\sum_{q}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}\Sigma\omega_{{\rm l},q}}}\frac{1}{\varepsilon+r_{\infty}q}\left(a_{\mathbf{q}}e^{i\mathbf{q}\cdot\mathbf{r}}-a_{\mathbf{q}}^{\dagger}e^{-i\mathbf{q}\cdot\mathbf{r}}\right).$ (S46)
IV Electron-polarons
IV.1 Hamiltonian and unitary transformations
The quantization of Eq. (S41) results in Hamiltonian for LO phonons, reading as
$\hat{H}_{\rm ph}=\sum_{k}\frac{\hbar\omega_{{\rm l},k}}{2}(\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}}^{\dagger}+\hat{a}_{-\mathbf{k}}^{\dagger}\hat{a}_{-\mathbf{k}})=\sum_{k}\hbar\omega_{{\rm l},k}(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+1/2)$ (S47)
The overall electron-LO phonon Hamiltonian is then
$\hat{H}_{\rm e-ph}=\sum_{k}\hbar\omega_{{\rm l},k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+\sum_{k}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}^{\dagger}\right)+\frac{\hat{\mathbf{p}}_{e}^{2}}{2m_{e}}$ (S48)
where
$V_{k}=ie\omega_{\rm t}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}\Sigma\omega_{{\rm l},k}}}\frac{1}{\varepsilon+r_{\infty}k}.$ (S49)
We further apply a variational approach for the calculation of polaron binding energy and charge carrier mass renormalization. In virtue of Ref. [6], we introduce the total momentum operator
$\hat{\mathbf{p}}=\sum_{k}\hbar\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+\hat{\mathbf{p}}_{e}.$ (S50)
One has
$[\hat{\mathbf{p}},\hat{\mathbf{p}}_{e}^{2}]=0,\qquad[\hat{\mathbf{p}},\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}]=0,\qquad[\hat{\mathbf{p}},\hat{a}_{\mathbf{q}}e^{i\mathbf{q}\cdot\mathbf{r}_{e}}]=\sum_{k}\hbar\mathbf{k}[\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}},\hat{a}_{\mathbf{q}}]e^{i\mathbf{q}\cdot\mathbf{r}_{e}}+\hat{a}_{\mathbf{q}}[\hat{\mathbf{p}},e^{i\mathbf{q}\cdot\mathbf{r}_{e}}]=0,$ (S51)
so that
$[\hat{\mathbf{p}},\hat{H}_{\rm e-ph}]=0,$ (S52)
ensuring the total momentum conservation. We then eliminate electron coordinates performing a unitary transformation with operator
$\hat{S}=e^{i(\mathbf{p}/\hbar-\sum_{k}\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}})\cdot{\mathbf{r}_{e}}}.$ (S53)
One has
$\tilde{\hat{\mathbf{p}}}_{e}$ $=\hat{\mathbf{p}}_{e}+\mathbf{p}-\hbar\sum_{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\mathbf{k},$ (S54)
so that
$\tilde{\hat{\mathbf{p}}}_{e}^{2}$ $=\hat{S}^{-1}\hat{\mathbf{p}}_{e}^{2}\hat{S}=\hat{S}^{-1}\hat{\mathbf{p}}_{e}\hat{S}\cdot\hat{S}^{-1}\hat{\mathbf{p}}_{e}\hat{S}=\left(\hat{\mathbf{p}}_{e}+\mathbf{p}-\hbar\sum_{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\mathbf{k}\right)^{2}.$ (S55)
Similarly,
$\tilde{\hat{a}}_{\mathbf{k}}$ $=\hat{a}_{\mathbf{k}}e^{-i\mathbf{k}\cdot\mathbf{r}_{e}},$ (S56)
meaning that
$\tilde{\hat{a}}_{\mathbf{k}}^{\dagger}$ $=\hat{a}_{\mathbf{k}}^{\dagger}e^{i\mathbf{k}\cdot\mathbf{r}_{e}},$
$\tilde{\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}}$ $=\hat{S}^{-1}\hat{a}_{\mathbf{k}}^{\dagger}\hat{S}\cdot\hat{S}^{-1}\hat{a}_{\mathbf{k}}\hat{S}=\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}},$
$\tilde{\hat{\mathbf{p}}}$ $=\hat{\mathbf{p}}_{e}+\mathbf{p}.$ (S57)
Now the transformed Hamiltonian is written as
$\tilde{\hat{H}}_{\rm e-ph}=\sum_{k}\hbar\omega_{{\rm l},k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+\sum_{k}\left(V_{k}\hat{a}_{\mathbf{k}}+V_{k}^{*}\hat{a}_{\mathbf{k}}^{\dagger}\right)+\frac{\left(\mathbf{p}-\sum_{k}\hbar\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\right)^{2}}{2m_{e}},$ (S58)
where we omit $\hat{\mathbf{p}}_{e}$ provided that we get rid of electron coordinate, so that $\hat{\mathbf{p}}_{e}(\hat{S}^{-1}\psi)=0$. We apply a variational anzats in a form $\psi=\hat{U}\psi_{0}$, where $\psi_{0}$ corresponds to phonon vacuum state $\hat{a}_{\mathbf{k}}\psi_{0}=0$, $\langle\psi_{0}|\psi_{0}\rangle=1$, and unitary operator
$\hat{U}=e^{\sum_{k}(\hat{a}_{\mathbf{k}}^{\dagger}f(\mathbf{k})-\hat{a}_{\mathbf{k}}f^{*}(\mathbf{k}))},$ (S59)
where $f(\mathbf{k})$ corresponds to a mean-field amplitude of phonon field minimizing the polaron energy. One has
$\hat{U}^{-1}\hat{a}_{\mathbf{k}}^{\dagger}\hat{U}$ $=\hat{a}_{\mathbf{k}}^{\dagger}-\sum_{q}[(\hat{a}_{\mathbf{q}}^{\dagger}f(\mathbf{q})-\hat{a}_{\mathbf{q}}f^{*}(\mathbf{q})),\hat{a}_{\mathbf{k}}^{\dagger}]+\ldots=\hat{a}_{\mathbf{k}}^{\dagger}-\sum_{q}(-1)\delta_{\mathbf{q},\mathbf{k}}f^{*}(\mathbf{q})=\hat{a}_{\mathbf{k}}^{\dagger}+f^{*}(\mathbf{k}),$ (S60)
and
$\hat{U}^{-1}\hat{a}_{\mathbf{k}}\hat{U}$ $=\hat{a}_{\mathbf{k}}+f(\mathbf{k}).$ (S61)
We evaluate the Hamiltonian terms one by one as
$\hat{U}^{-1}\sum_{k}\hbar\omega_{{\rm l},k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{U}$ $=\sum_{k}\hbar\omega_{{\rm l},k}\left(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+f(\mathbf{k})\hat{a}_{\mathbf{k}}^{\dagger}+f^{*}(\mathbf{k})\hat{a}_{\mathbf{k}}+f^{*}(\mathbf{k})f(\mathbf{k})\right)$ (S62)
$\hat{U}^{-1}\sum_{k}\left(V_{k}\hat{a}_{\mathbf{k}}+V_{k}^{*}\hat{a}_{\mathbf{k}}^{\dagger}\right)\hat{U}$ $=\sum_{k}\left(V_{k}\hat{a}_{\mathbf{k}}+V_{k}f(\mathbf{k})+V_{k}^{*}\hat{a}_{\mathbf{k}}^{\dagger}+V_{k}^{*}f^{*}(\mathbf{k})\right)$ (S63)
$\hat{U}^{-1}\frac{\mathbf{p}^{2}}{2m_{e}}\hat{U}$ $=\frac{\mathbf{p}^{2}}{2m_{e}}$ (S64)
$-\frac{1}{m}\hat{U}^{-1}\mathbf{p}\sum_{k}\hbar\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{U}$ $=-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}\left(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+f(\mathbf{k})\hat{a}_{\mathbf{k}}^{\dagger}+f^{*}(\mathbf{k})\hat{a}_{\mathbf{k}}+f^{*}(\mathbf{k})f(\mathbf{k})\right)$ (S65)
and
$\hat{U}^{-1}\frac{1}{2m_{e}}\sum_{k}\sum_{k^{\prime}}\hbar\mathbf{k}\cdot\hbar\mathbf{k}^{\prime}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}\hat{U}=\frac{\hbar^{2}}{2m_{e}}\sum_{k}\sum_{k^{\prime}}\mathbf{k}\cdot\mathbf{k}^{\prime}(\hat{a}_{\mathbf{k}}^{\dagger}+f^{*}(\mathbf{k}))(\hat{a}_{\mathbf{k}}+f(\mathbf{k}))(\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}+f^{*}(\mathbf{k}^{\prime}))(\hat{a}_{\mathbf{k}^{\prime}}+f(\mathbf{k}^{\prime}))$
$=\frac{\hbar^{2}}{2m_{e}}\sum_{k}\sum_{k^{\prime}}\mathbf{k}\cdot\mathbf{k}^{\prime}\left(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}}f(\mathbf{k}^{\prime})+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k}^{\prime})+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}|f(\mathbf{k}^{\prime})|^{2}\right.$
$\hskip 100.0pt+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}f(\mathbf{k})+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}f(\mathbf{k}^{\prime})f(\mathbf{k})+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k}^{\prime})f(\mathbf{k})+\hat{a}_{\mathbf{k}}^{\dagger}|f(\mathbf{k}^{\prime})|^{2}f(\mathbf{k})$
$\hskip 100.0pt+\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k})+\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}}f(\mathbf{k}^{\prime})f^{*}(\mathbf{k})+\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k}^{\prime})f^{*}(\mathbf{k})+\hat{a}_{\mathbf{k}}|f(\mathbf{k}^{\prime})|^{2}f^{*}(\mathbf{k})$
$\hskip 100.0pt\left.+\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}|f(\mathbf{k})|^{2}+\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}f(\mathbf{k}^{\prime})|f(\mathbf{k})|^{2}+\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k}^{\prime})|f(\mathbf{k})|^{2}+|f(\mathbf{k})|^{2}|f(\mathbf{k}^{\prime})|^{2}\right)$
$+\frac{\hbar^{2}}{2m_{e}}\sum_{k}\mathbf{k}^{2}\left(\hat{a}_{\mathbf{k}}^{\dagger}f(\mathbf{k})+\hat{a}_{\mathbf{k}}f^{*}(\mathbf{k})+|f(\mathbf{k})|^{2}\right)$ (S66)
Now the Hamiltonian can be presented in the form ordered by the number of phonon creation / annihilation operators:
$\hat{H}_{\rm e-ph}=\hat{H}_{0}+\hat{H}_{1}+\hat{H}_{2}+\hat{H}_{3}+\hat{H}_{4},$ (S67)
where
$\hat{H}_{0}$ $=\frac{\mathbf{p}^{2}}{2m_{e}}+\sum_{k}\left(V_{k}f(\mathbf{k})+V_{k}^{*}f^{*}(\mathbf{k})\right)+\frac{\hbar^{2}}{2m_{e}}\left(\sum_{k}|f(\mathbf{k})|^{2}\mathbf{k}\right)^{2}+\sum_{k}|f(\mathbf{k})|^{2}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)$ (S68)
$\hat{H}_{1}$ $=\sum_{k}\left[V_{k}+f^{*}(\mathbf{k})\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+\frac{\hbar^{2}}{m_{e}}\mathbf{k}\cdot\left(\sum_{k^{\prime}}|f(\mathbf{k}^{\prime})|^{2}\mathbf{k}^{\prime}\right)\right)\right]\hat{a}_{\mathbf{k}}$
$+\sum_{k}\left[V_{k}^{*}+f(\mathbf{k})\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+\frac{\hbar^{2}}{m_{e}}\mathbf{k}\cdot\left(\sum_{k^{\prime}}|f(\mathbf{k}^{\prime})|^{2}\mathbf{k}^{\prime}\right)\right)\right]\hat{a}_{\mathbf{k}}^{\dagger}$ (S69)
$\hat{H}_{2}$ $=\sum_{k}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}}{m_{e}}\mathbf{k}\cdot\left(\sum_{k^{\prime}}\mathbf{k}^{\prime}|f(\mathbf{k}^{\prime})|^{2}\right)\right)\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}$ (S70)
$+\frac{\hbar^{2}}{2m_{e}}\sum_{k}\sum_{k^{\prime}}\mathbf{k}\cdot\mathbf{k}^{\prime}\left(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}f(\mathbf{k}^{\prime})f(\mathbf{k})+\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k}^{\prime})f^{*}(\mathbf{k})+2\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k}^{\prime})f(\mathbf{k})\right)$ (S71)
$\hat{H}_{3}$ $=\frac{\hbar^{2}}{m_{e}}\sum_{k}\sum_{k^{\prime}}\mathbf{k}\cdot\mathbf{k}^{\prime}\left(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{k}^{\prime}}f^{*}(\mathbf{k}^{\prime})+\hat{a}_{\mathbf{k}^{\prime}}^{\dagger}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}f(\mathbf{k}^{\prime})\right)$ (S72)
$\hat{H}_{4}$ $=\frac{\left(\sum_{k}\hbar\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\right)^{2}}{2m_{e}}$ (S73)
IV.2 Polaron energy minimization
Taking into account that $\hat{a}_{\mathbf{k}}\psi_{0}=0$, we calculate the variational energy as
$E$ $=\langle\psi|\hat{H}_{\rm e-ph}|\psi\rangle=\langle\psi|\hat{H}_{0}|\psi\rangle$
$=\frac{\mathbf{p}^{2}}{2m_{e}}+\sum_{k}\left(V_{k}f(\mathbf{k})+V_{k}^{*}f^{*}(\mathbf{k})\right)+\frac{\hbar^{2}}{2m_{e}}\left(\sum_{k}|f(\mathbf{k})|^{2}\mathbf{k}\right)^{2}+\sum_{k}|f(\mathbf{k})|^{2}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right),$ (S74)
which can be minimized setting
$\frac{\delta E}{\delta f(\mathbf{k})}=\frac{\delta E}{\delta f(\mathbf{k}^{\prime})}=0,$ (S75)
resulting in
$V_{k}+f^{*}(\mathbf{k})\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+\frac{\hbar^{2}}{m_{e}}\mathbf{k}\cdot\left(\sum_{k^{\prime}}|f(\mathbf{k}^{\prime})|^{2}\mathbf{k}^{\prime}\right)\right)=0,$
$V_{k}^{*}+f(\mathbf{k})\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+\frac{\hbar^{2}}{m_{e}}\mathbf{k}\cdot\left(\sum_{k^{\prime}}|f(\mathbf{k}^{\prime})|^{2}\mathbf{k}^{\prime}\right)\right)=0.$ (S76)
We note that the obtained expressions correspond to the sum coefficients in $\hat{H}_{1}$, meaning the exact cancellation of linear in $\hat{a}_{\mathbf{k}}$ terms from the Hamiltonian. Noting that the only preferred direction is that of $\mathbf{p}$, one may write
$\beta\mathbf{p}=\sum_{k}|f(\mathbf{k})|^{2}\hbar\mathbf{k},$ (S77)
and obtain an equation for amplitude $f(\mathbf{k})$:
$f(\mathbf{k})=-\frac{V_{k}^{*}}{\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)+\frac{\hbar^{2}k^{2}}{2m_{e}}}.$ (S78)
Plugging in, we get
$\beta\mathbf{p}=\sum_{k}|f(\mathbf{k})|^{2}\hbar\mathbf{k}=\sum_{k}\frac{|V_{k}|^{2}\hbar\mathbf{k}}{\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}.$ (S79)
As for energy, we note that
$\frac{\hbar^{2}}{2m_{e}}\left(\sum_{k}|f(\mathbf{k})|^{2}\mathbf{k}\right)^{2}$ $=\frac{\hbar^{2}}{m_{e}}\left(\sum_{k}|f(\mathbf{k})|^{2}\mathbf{k}\right)^{2}-\frac{\hbar^{2}}{2m_{e}}\left(\sum_{k}|f(\mathbf{k})|^{2}\mathbf{k}\right)^{2}=\frac{\hbar}{m_{e}}\beta\mathbf{p}\cdot\sum_{k}|f(\mathbf{k})|^{2}\mathbf{k}-\frac{1}{2m_{e}}(\beta\mathbf{p})^{2}$ (S80)
resulting in
$E$ $=\frac{\mathbf{p}^{2}}{2m_{e}}(1-\beta^{2})-\sum_{k}\frac{2|V_{k}|^{2}}{\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\sum_{k}\frac{\hbar}{m_{e}}\beta\mathbf{p}\cdot\mathbf{k}\frac{|V_{k}|^{2}}{\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}$
$+\sum_{k}\frac{|V_{k}|^{2}}{\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)$
$=\frac{\mathbf{p}^{2}}{2m_{e}}(1-\beta^{2})-\sum_{k}\frac{|V_{k}|^{2}}{\hbar\omega_{{\rm l},k}-\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)+\frac{\hbar^{2}k^{2}}{2m_{e}}}.$ (S81)
Next, one has
$|V_{k}|^{2}$ $=\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{\Sigma(4\varepsilon_{0})}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)}}\frac{1}{(\varepsilon+r_{\infty}k)}.$ (S82)
Expanding the Eq. (S79) up to first order in $\mathbf{k}$, we get
$\beta\mathbf{p}=\sum_{k}\frac{|V_{k}|^{2}\hbar\mathbf{k}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{3}}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+2\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)\right),$ (S83)
Plugging in the values of $|V_{k}|^{2}$, $\omega_{{\rm l},k}$, and proceeding with integration, we reach at
$\beta p$ $=\frac{\Sigma}{(2\pi)^{2}}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{\Sigma(4\varepsilon_{0})}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+2\frac{\hbar}{m_{e}}pk\cos\varphi(1-\beta)\right)\frac{\hbar k\cos\varphi k{\rm d}k{\rm d}\varphi}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{3}}$
$=\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{(8\pi\varepsilon_{0})}\int\limits_{0}^{\infty}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\frac{\hbar}{m_{e}}pk(1-\beta)\frac{\hbar kk{\rm d}k}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{3}}$
$=\frac{e^{2}(r_{0}-r_{\infty})}{(8\pi\varepsilon_{0})}\frac{p(1-\beta)}{m_{e}\omega_{\rm t}^{2}}\int\limits_{0}^{\infty}\frac{1}{(\varepsilon+r_{0}k)^{2}}\frac{k^{3}{\rm d}k}{\left(1+r_{e,{\rm t}}^{2}k^{2}\sqrt{\frac{\varepsilon+r_{\infty}k}{\varepsilon+r_{0}k}}\right)^{3}},$ (S84)
Similarly
$E$ $\approx\frac{p^{2}}{2m_{e}}(1-\beta^{2})-\sum_{k}\frac{|V_{k}|^{2}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\left(1+\frac{\frac{\hbar}{m_{e}}\mathbf{p}\cdot\mathbf{k}(1-\beta)}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{\frac{\hbar^{2}}{m_{e}^{2}}(1-\beta)^{2}(\mathbf{p}\cdot\mathbf{k})^{2}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}\right)$
$=\frac{p^{2}}{2m_{e}}(1-\beta^{2})-\frac{\Sigma}{(2\pi)^{2}}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{|V_{k}|^{2}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\left(1+\frac{\frac{\hbar}{m_{e}}pk\cos\varphi(1-\beta)}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{\frac{\hbar^{2}}{m_{e}^{2}}(1-\beta)^{2}(pk\cos\varphi)^{2}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}\right)k{\rm d}k{\rm d}\varphi$
$=\frac{p^{2}}{2m_{e}}(1-\beta^{2})-\int\limits_{0}^{\infty}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{(8\pi\varepsilon_{0})}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)}}\frac{1}{(\varepsilon+r_{\infty}k)}\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\left(1+\frac{\frac{\hbar^{2}}{2m_{e}^{2}}(1-\beta)^{2}(pk)^{2}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}\right)k{\rm d}k$
The first term in integral corresponds to polaron binding energy, and has a form
$E_{\rm b}$ $=-\frac{e^{2}(r_{0}-r_{\infty})}{8\pi\varepsilon_{0}}\int\limits_{0}^{\infty}\frac{1}{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)}\frac{k{\rm d}k}{1+r_{e,{\rm t}}^{2}k^{2}\sqrt{\frac{\varepsilon+r_{\infty}k}{\varepsilon+r_{0}k}}}.$ (S86)
Similarly, the second term has a form
$E_{2}=-\frac{e^{2}(r_{0}-r_{\infty})}{(8\pi\varepsilon_{0})(\hbar\omega_{\rm t})^{2}}\frac{\hbar^{2}}{2m_{e}^{2}}(1-\beta)^{2}p^{2}\int\limits_{0}^{\infty}\frac{1}{(\varepsilon+r_{0}k)^{2}}\frac{k^{3}{\rm d}k}{\left(1+\frac{\hbar k^{2}}{2\omega_{\rm t}m_{e}}\sqrt{\frac{\varepsilon+r_{\infty}k}{\varepsilon+r_{0}k}}\right)^{3}}$
Introducing dimensionless quantities
$r_{\infty}k/\varepsilon=\tilde{k},\quad r_{0}/r_{\infty}=\sigma_{0},\quad\varepsilon r_{e,{\rm t}}/r_{\infty}=\sigma_{e,{\rm t}},$ (S88)
we can rewrite
$E$ $=-\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\int\limits_{0}^{\infty}\frac{1}{(1+\sigma_{0}\tilde{k})(1+\tilde{k})}\frac{\tilde{k}{\rm d}\tilde{k}}{1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}$
$+\frac{p^{2}}{2m_{e}}\left[(1-\beta^{2})-\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\frac{1}{(r_{\infty}\hbar\omega_{\rm t})^{2}}\frac{\hbar^{2}}{m_{e}}(1-\beta)^{2}\int\limits_{0}^{\infty}\frac{1}{(1+\sigma_{0}\tilde{k})^{2}}\frac{\tilde{k}^{3}{\rm d}\tilde{k}}{\left(1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}\right)^{3}}\right]$
$=-E_{0}I_{\rm b}+\frac{p^{2}}{2m_{e}}\left[(1-\beta^{2})-\frac{E_{0}}{\hbar\omega_{\rm t}}I_{m}(1-\beta)^{2}\right],$ (S89)
where
$E_{0}=\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}},$ (S90)
$I_{\rm b}$ $=\int\limits_{0}^{\infty}\frac{\sigma_{0}-1}{2(1+\sigma_{0}\tilde{k})(1+\tilde{k})}\frac{\tilde{k}{\rm d}\tilde{k}}{1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}},$ (S91)
$I_{m}$ $=\int\limits_{0}^{\infty}\frac{\sigma_{e,{\rm t}}^{2}(\sigma_{0}-1)}{(1+\sigma_{0}\tilde{k})^{2}}\frac{\tilde{k}^{3}{\rm d}\tilde{k}}{\left(1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}\right)^{3}}.$ (S92)
Similarly, we rewrite
$\beta$ $=\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\frac{(1-\beta)}{m_{e}\omega_{\rm t}^{2}r_{\infty}^{2}}\int\limits_{0}^{\infty}\frac{1}{(1+\sigma_{0}\tilde{k})^{2}}\frac{\tilde{k}^{3}{\rm d}\tilde{k}}{\left(1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}\right)^{3}}=\frac{E_{0}}{\hbar\omega_{\rm t}}I_{m}(1-\beta).$ (S93)
Denoting
$\alpha_{e}=\frac{E_{0}}{\hbar\omega_{\rm t}}I_{m}(\sigma_{0},\sigma_{e,{\rm t}}),$ (S94)
we get
$\beta=\frac{\alpha_{e}}{1+\alpha_{e}},$ (S95)
and
$E$ $=-E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})+\frac{p^{2}}{2m_{e}}\frac{1}{1+\alpha_{e}}.$ (S96)
We can further proceed with approximate expressions for the integral $I_{rmb}$. Applying an approximation $\sqrt{(1+\tilde{k})/(1+\sigma_{0}\tilde{k})}\approx 1/\sqrt{\sigma_{0}}$, we get
$I_{\rm b}\approx\frac{\sigma_{0}-1}{2(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})}\left(\frac{\pi(\sigma_{0}+1)\tilde{\sigma}_{e,{\rm t}}}{2\left(\tilde{\sigma}_{e,{\rm t}}^{2}+1\right)}+\frac{\left(\tilde{\sigma}_{e,{\rm t}}^{2}-\sigma_{0}\right)\log(\tilde{\sigma_{e,{\rm t}}})}{\tilde{\sigma}_{e,{\rm t}}^{2}+1}-\frac{\sigma_{0}\log(\sigma_{0})}{\sigma_{0}-1}\right),$ (S97)
where $\tilde{\sigma}_{e,{\rm t}}=\sigma_{e,{\rm t}}/\sigma_{0}^{1/4}$. Similarly,
$I_{m}\approx\frac{\sigma_{e,{\rm t}}^{2}(\sigma_{0}-1)}{8\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})^{4}}\left[2(\sigma_{0}^{6}+\tilde{\sigma}_{e,{\rm t}}^{6})-3\pi\sigma_{0}\tilde{\sigma}_{e,{\rm t}}\left(\sigma_{0}^{4}+\tilde{\sigma}_{e,{\rm t}}^{4}\right)-18\sigma_{0}^{2}\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2}-\pi)+24\sigma_{0}^{2}\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}-\tilde{\sigma}_{e,{\rm t}}^{2})\log\!\biggl{(}\sigma_{0}/\tilde{\sigma}_{e,{\rm t}}\biggr{)}\right].$ (S98)
V Exciton states
We next consider the exciton-polaron problem. The Hamiltonian of an electron-hole pair interacting with LO phonons can be written as
$\hat{H}_{\rm X}$ $=\frac{\hat{\mathbf{p}}_{e}^{2}}{2m_{e}}+\frac{\hat{\mathbf{p}}_{h}^{2}}{2m_{h}}+V_{\rm KR}(|\mathbf{r}_{e}-\mathbf{r}_{h}|,r_{\infty})+\sum_{k}\hbar\omega_{{\rm l},k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}$
$+\sum_{k}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{e}}\hat{a}_{\mathbf{k}}^{\dagger}\right)-\sum_{k}\left(V_{k}e^{i\mathbf{k}\cdot\mathbf{r}_{h}}\hat{a}_{\mathbf{k}}+V_{k}^{*}e^{-i\mathbf{k}\cdot\mathbf{r}_{h}}\hat{a}_{\mathbf{k}}^{\dagger}\right),$ (S99)
where $V_{\rm KR}$ is the bare 2D Coulomb potential. Proceeding to the center-of-mass and relative coordinates, one can rewrite
$\hat{H}_{\rm X}=\frac{\hat{\mathbf{P}}^{2}}{2M}+\frac{\hat{\mathbf{p}}^{2}}{2\mu}+V_{\rm KR}(r,r_{\infty})+\sum_{k}\hbar\omega_{{\rm l},k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+\sum_{k}\left(V_{k}\rho_{\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{R}}\hat{a}_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}e^{-i\mathbf{k}\cdot\mathbf{R}}\hat{a}_{\mathbf{k}}^{\dagger}\right),$ (S100)
where
$\rho_{\mathbf{k}}=e^{im_{h}\mathbf{k}\cdot\mathbf{r}/M}-e^{-im_{e}\mathbf{k}\cdot\mathbf{r}/M}.$ (S101)
We perform an unitary transform with the following operator
$\hat{U}_{1}=e^{i(\mathbf{Q}-\sum_{k}\mathbf{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}})\cdot\mathbf{R}},$ (S102)
where $\hbar\mathbf{Q}$ corresponds to overall center-of-mass momentum. Similarly to the case of single polaron, we get the following relations
$\hat{U}_{1}^{-1}\hat{\mathbf{P}}\hat{U}_{1}$ $=\hat{\mathbf{P}}+\hbar\mathbf{Q}-\hbar\sum_{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\mathbf{k},$ (S103)
$\hat{U}_{1}^{-1}\hat{\mathbf{p}}\hat{U}_{1}$ $=\hat{\mathbf{p}},$ (S104)
$\hat{U}_{1}^{-1}\hat{a}_{\mathbf{k}}\hat{U}_{1}$ $=\hat{a}_{\mathbf{k}}e^{-i\mathbf{k}\cdot\mathbf{R}},$ (S105)
yielding in the transformed Hamiltonian
$\tilde{\hat{H}}_{\rm X}=\frac{(\hat{\mathbf{P}}+\hbar\mathbf{Q}-\hbar\sum_{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\mathbf{k})^{2}}{2M}+\frac{\hat{\mathbf{p}}^{2}}{2\mu}+V_{\rm KR}(r,r_{\infty})+\sum_{k}\hbar\omega_{{\rm l},k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+\sum_{k}\left(V_{k}\rho_{\mathbf{k}}\hat{a}_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}\hat{a}_{\mathbf{k}}^{\dagger}\right).$ (S106)
We get rid of $\mathbf{R}$-coordinate in Hamiltonian, so that we may set $\hat{\mathbf{P}}+\hbar\mathbf{Q}\rightarrow\hbar\mathbf{Q}$. Expanding the square term of Hamiltonian, we get
$\tilde{\hat{H}}_{\rm X}$ $=\frac{\hbar^{2}\mathbf{Q}^{2}}{2M}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\sum_{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\mathbf{k}+\frac{\hbar^{2}}{2M}\left(\sum_{k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\mathbf{k}\right)^{2}+\frac{\hat{\mathbf{p}}^{2}}{2\mu}+V_{\rm KR}(r,r_{\infty})+\sum_{k}\hbar\omega_{{\rm l},k}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+\sum_{k}\left(V_{k}\rho_{\mathbf{k}}\hat{a}_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}\hat{a}_{\mathbf{k}}^{\dagger}\right)$
$=\frac{\hbar^{2}\mathbf{Q}^{2}}{2M}+\frac{\hat{\mathbf{p}}^{2}}{2\mu}+V_{\rm KR}(r,r_{\infty})+\sum_{k}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2M}\right)\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+\frac{\hbar^{2}}{2M}\sum_{k}\sum_{q}\mathbf{k}\cdot\mathbf{q}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}\hat{a}_{\mathbf{k}}$
$+\sum_{k}\left(V_{k}\rho_{\mathbf{k}}\hat{a}_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}\hat{a}_{\mathbf{k}}^{\dagger}\right).$ (S107)
We proceed with another transformation of Hamiltonian with unitary operator
$\hat{U}_{2}=e^{\sum_{k}(\hat{a}_{\mathbf{k}}F^{*}_{\mathbf{k}}(\mathbf{r})-\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{k}}(\mathbf{r}))}$ (S108)
where $F_{\mathbf{k}}(\mathbf{r})$ being the mean field value of the displacement. We then have
$\hat{U}_{2}^{-1}\hat{a}_{\mathbf{k}}\hat{U}_{2}=\hat{a}_{\mathbf{k}}+\left[\sum_{q}(-\hat{a}_{\mathbf{q}}F^{*}_{\mathbf{q}}(\mathbf{r})+\hat{a}_{\mathbf{q}}^{\dagger}F_{\mathbf{q}}(\mathbf{r})),\hat{a}_{\mathbf{k}}\right]+\cdots=\hat{a}_{\mathbf{k}}-\sum_{q}\delta_{\mathbf{q},\mathbf{k}}F_{\mathbf{q}}(\mathbf{r})=\hat{a}_{\mathbf{k}}-F_{\mathbf{k}}(\mathbf{r}),$ (S109)
$\hat{U}_{2}^{-1}\hat{a}_{\mathbf{k}}^{\dagger}\hat{U}_{2}=\hat{a}_{\mathbf{k}}^{\dagger}-F_{\mathbf{k}}^{*}(\mathbf{r}),$ (S110)
and
$\hat{U}_{2}^{-1}\hat{\mathbf{p}}\hat{U}_{2}=\hat{\mathbf{p}}+\left[\sum_{k}(-\hat{a}_{\mathbf{k}}F^{*}_{\mathbf{k}}(\mathbf{r})+\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{k}}(\mathbf{r})),\hat{\mathbf{p}}\right]+\frac{1}{2}\left[\sum_{q}(-\hat{a}_{\mathbf{q}}F^{*}_{\mathbf{q}}(\mathbf{r})+\hat{a}_{\mathbf{q}}^{\dagger}F_{\mathbf{q}}(\mathbf{r})),\left[\sum_{k}(-\hat{a}_{\mathbf{k}}F^{*}_{\mathbf{k}}(\mathbf{r})+\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{k}}(\mathbf{r})),\hat{\mathbf{p}}\right]\right]+\cdots$
$=\hat{\mathbf{p}}-2\mu\frac{\hbar}{2i\mu}\sum_{k}\left(\hat{a}_{\mathbf{k}}^{\dagger}\nabla F_{\mathbf{k}}(\mathbf{r})-\hat{a}_{\mathbf{k}}\nabla F^{*}_{\mathbf{k}}(\mathbf{r})\right)+\mu\frac{\hbar}{2i\mu}\sum_{k}\left(F^{*}_{\mathbf{k}}(\mathbf{r})\nabla F_{\mathbf{k}}(\mathbf{r})-F_{\mathbf{k}}(\mathbf{r})\nabla F^{*}_{\mathbf{k}}(\mathbf{r})\right)=\hat{\mathbf{p}}-2\mu\mathbf{J}(\mathbf{r},\hat{a}_{\mathbf{k}})+\mu\mathbf{j}(\mathbf{r}).$ (S111)
The Hamiltonian terms are transformed as follows:
$\hat{U}_{2}^{-1}\left(\frac{\hat{\mathbf{p}}^{2}}{2\mu}\right)\hat{U}_{2}=\frac{1}{2\mu}\hat{U}_{2}^{-1}\hat{\mathbf{p}}\hat{U}_{2}\cdot\hat{U}_{2}^{-1}\hat{\mathbf{p}}\hat{U}_{2}=\frac{1}{2\mu}\left(\hat{\mathbf{p}}-2\mu\mathbf{J}(\mathbf{r},\hat{a}_{\mathbf{k}})+\mu\mathbf{j}(\mathbf{r})\right)^{2}$
$=\frac{1}{2\mu}\left(\hat{\mathbf{p}}^{2}+\mu^{2}\mathbf{j}^{2}+\mu(\hat{\mathbf{p}}\cdot\mathbf{j}+\mathbf{j}\cdot\hat{\mathbf{p}})+4\mu^{2}\mathbf{J}^{2}-4\mu^{2}\mathbf{J}\cdot\mathbf{j}-2\mu(\hat{\mathbf{p}}\cdot\mathbf{J}+\mathbf{J}\cdot\hat{\mathbf{p}})\right)$
$=\frac{\left(\hat{\mathbf{p}}+\mu\mathbf{j}\right)^{2}}{2\mu}+\frac{\hbar^{2}}{2\mu}\sum_{k}\nabla F^{*}_{\mathbf{k}}\cdot\nabla F_{\mathbf{k}}+\frac{\hbar^{2}}{2\mu}\sum_{k}\sum_{q}\left(2\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}\nabla F_{\mathbf{k}}\cdot\nabla F^{*}_{\mathbf{q}}-\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}\nabla F_{\mathbf{k}}\cdot\nabla F_{\mathbf{q}}-\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{q}}\nabla F^{*}_{\mathbf{k}}\cdot\nabla F^{*}_{\mathbf{q}}\right)$
$+\frac{\hbar}{i\mu}\left(\hat{a}_{\mathbf{k}}\nabla F^{*}_{\mathbf{k}}-\hat{a}_{\mathbf{k}}^{\dagger}\nabla F_{\mathbf{k}}\right)\cdot\left(\hat{\mathbf{p}}+\mu\mathbf{j}\right)+\frac{\hbar^{2}}{2\mu}\left(\hat{a}_{\mathbf{k}}^{\dagger}\Delta F_{\mathbf{k}}-\hat{a}_{\mathbf{k}}\Delta F^{*}_{\mathbf{k}}\right),$ (S112)
$\hat{U}_{2}^{-1}\left(\frac{\hbar^{2}\mathbf{Q}^{2}}{2M}+V_{\rm KR}(r,r_{\infty})\right)\hat{U}_{2}$ $=\frac{\hbar^{2}\mathbf{Q}^{2}}{2M}+V_{\rm KR}(r,r_{\infty}),$ (S113)
$\hat{U}_{2}^{-1}\left(\sum_{k}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2M}\right)\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\right)\hat{U}_{2}$ $=\sum_{k}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2M}\right)\left(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}+F_{\mathbf{k}}^{*}F_{\mathbf{k}}-\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{k}}-\hat{a}_{\mathbf{k}}F_{\mathbf{k}}^{*}\right),$ (S114)
$\hat{U}_{2}^{-1}\left(\sum_{k}\left(V_{k}\rho_{\mathbf{k}}\hat{a}_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}\hat{a}_{\mathbf{k}}^{\dagger}\right)\right)\hat{U}_{2}$ $=\sum_{k}\left(V_{k}\rho_{\mathbf{k}}\hat{a}_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}\hat{a}_{\mathbf{k}}^{\dagger}-V_{k}\rho_{\mathbf{k}}F_{\mathbf{k}}-V_{k}^{*}\rho_{\mathbf{k}}^{*}F_{\mathbf{k}}^{*}\right),$ (S115)
$\hat{U}_{2}^{-1}\left(\frac{\hbar^{2}}{2M}\sum_{k}\sum_{q}\mathbf{k}\cdot\mathbf{q}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}\hat{a}_{\mathbf{k}}\right)\hat{U}_{2}=\frac{\hbar^{2}}{2M}\sum_{k}\sum_{q}\mathbf{k}\cdot\mathbf{q}\left(\hat{a}_{\mathbf{k}}^{\dagger}-F_{\mathbf{k}}^{*}\right)\left(\hat{a}_{\mathbf{q}}^{\dagger}-F_{\mathbf{q}}^{*}\right)\left(\hat{a}_{\mathbf{q}}-F_{\mathbf{q}}\right)\left(\hat{a}_{\mathbf{k}}-F_{\mathbf{k}}\right)$
$=\frac{\hbar^{2}}{2M}\sum_{k}\sum_{q}\mathbf{k}\cdot\mathbf{q}\left(\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}\hat{a}_{\mathbf{k}}-\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}F_{\mathbf{k}}-\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{F}_{\mathbf{q}}+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}F_{\mathbf{q}}F_{\mathbf{k}}\right.$
$\hskip 80.0pt-\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{q}}^{*}\hat{a}_{\mathbf{q}}\hat{a}_{\mathbf{k}}+\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{q}}^{*}\hat{a}_{\mathbf{q}}F_{\mathbf{k}}+\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{q}}^{*}\hat{a}_{\mathbf{k}}\hat{F}_{\mathbf{q}}-\hat{a}_{\mathbf{k}}^{\dagger}F_{\mathbf{q}}^{*}F_{\mathbf{q}}F_{\mathbf{k}}$
$\hskip 80.0pt-F_{\mathbf{k}}^{*}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}\hat{a}_{\mathbf{k}}+F_{\mathbf{k}}^{*}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}F_{\mathbf{k}}+F_{\mathbf{k}}^{*}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{F}_{\mathbf{q}}-F_{\mathbf{k}}^{*}\hat{a}_{\mathbf{q}}^{\dagger}F_{\mathbf{q}}F_{\mathbf{k}}$
$\hskip 80.0pt\left.+F_{\mathbf{k}}^{*}F_{\mathbf{q}}^{*}\hat{a}_{\mathbf{q}}\hat{a}_{\mathbf{k}}-F_{\mathbf{k}}^{*}F_{\mathbf{q}}^{*}\hat{a}_{\mathbf{q}}F_{\mathbf{k}}-F_{\mathbf{k}}^{*}F_{\mathbf{q}}^{*}\hat{a}_{\mathbf{k}}\hat{F}_{\mathbf{q}}+F_{\mathbf{k}}^{*}F_{\mathbf{q}}^{*}F_{\mathbf{q}}F_{\mathbf{k}}\right).$ (S116)
The transformed Hamiltonian then is
$\hat{H}_{\rm X}=\hat{H}_{\rm X0}+\hat{H}_{\rm X1}+\hat{H}_{\rm X2}+\hat{H}_{\rm X3}+\hat{H}_{\rm X4},$ (S117)
where
$\hat{H}_{\rm X0}$ $=\frac{\hbar^{2}\mathbf{Q}^{2}}{2M}+V_{\rm KR}(r,r_{\infty})+\frac{\left(\hat{\mathbf{p}}+\mu\mathbf{j}\right)^{2}}{2\mu}+\frac{\hbar^{2}}{2\mu}\sum_{k}\nabla F^{*}_{\mathbf{k}}\cdot\nabla F_{\mathbf{k}}+\sum_{k}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2M}\right)|F_{\mathbf{k}}|^{2}$
$-\sum_{k}\left(V_{k}\rho_{\mathbf{k}}F_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}F_{\mathbf{k}}^{*}\right)+\frac{\hbar^{2}}{2M}\sum_{k}\sum_{q}\mathbf{k}\cdot\mathbf{q}|F_{\mathbf{k}}|^{2}|F_{\mathbf{q}}|^{2}$ (S118)
$\hat{H}_{\rm X1}$ $=\sum_{k}\left(V_{k}\rho_{\mathbf{k}}-\left(\hbar\omega_{{\rm l},k}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2M}+\frac{\hbar^{2}}{M}\sum_{q}\mathbf{k}\cdot\mathbf{q}|F_{\mathbf{q}}|^{2}\right)F_{\mathbf{k}}^{*}-\frac{\hbar^{2}}{2\mu}\Delta F^{*}_{\mathbf{k}}+\frac{\hbar}{i\mu}\nabla F_{\mathbf{k}}^{*}\cdot\left(\hat{\mathbf{p}}+\mu\mathbf{j}\right)\right)\hat{a}_{\mathbf{k}}$
$+\sum_{k}\left(V_{k}^{*}\rho_{\mathbf{k}}^{*}-\left(\hbar\omega_{{\rm l},k}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2M}+\frac{\hbar^{2}}{M}\sum_{q}\mathbf{k}\cdot\mathbf{q}|F_{\mathbf{q}}|^{2}\right)F_{\mathbf{k}}+\frac{\hbar^{2}}{2\mu}\Delta F_{\mathbf{k}}-\frac{\hbar}{i\mu}\nabla F_{\mathbf{k}}\cdot\left(\hat{\mathbf{p}}+\mu\mathbf{j}\right)\right)\hat{a}_{\mathbf{k}}^{\dagger}$
$\hat{H}_{\rm X2}$ $=\sum_{k}\left(\hbar\omega_{{\rm l},k}-\frac{\hbar^{2}}{M}\mathbf{Q}\cdot\mathbf{k}+\frac{\hbar^{2}k^{2}}{2M}+\frac{\hbar^{2}}{M}\sum_{q}\mathbf{k}\cdot\mathbf{q}|F_{\mathbf{q}}|^{2}\right)\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}$
$+\sum_{k}\sum_{q}\left(\frac{\hbar^{2}}{M}\mathbf{k}\cdot\mathbf{q}F_{\mathbf{q}}^{*}F_{\mathbf{k}}+\frac{\hbar^{2}}{\mu}\nabla F_{\mathbf{k}}\cdot\nabla F^{*}_{\mathbf{q}}\right)\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}$
$+\sum_{k}\sum_{q}\left(\frac{\hbar^{2}}{2M}\mathbf{k}\cdot\mathbf{q}F_{\mathbf{q}}^{*}F_{\mathbf{k}}^{*}-\frac{\hbar^{2}}{2\mu}\nabla F_{\mathbf{k}}^{*}\cdot\nabla F^{*}_{\mathbf{q}}\right)\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{q}}$
$+\sum_{k}\sum_{q}\left(\frac{\hbar^{2}}{2M}\mathbf{k}\cdot\mathbf{q}F_{\mathbf{q}}F_{\mathbf{k}}-\frac{\hbar^{2}}{2\mu}\nabla F_{\mathbf{k}}\cdot\nabla F_{\mathbf{q}}\right)\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}$ (S119)
$\hat{H}_{\rm X3}$ $=-\frac{\hbar^{2}}{M}\sum_{k}\sum_{q}\mathbf{k}\cdot\mathbf{q}\left(\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{F}_{\mathbf{q}}+\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{q}}\hat{F}_{\mathbf{q}}^{*}\right)$ (S120)
$\hat{H}_{\rm X4}$ $=\frac{\hbar^{2}}{2M}\sum_{k}\sum_{q}\mathbf{k}\cdot\mathbf{q}\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{q}}^{\dagger}\hat{a}_{\mathbf{q}}\hat{a}_{\mathbf{k}}$ (S121)
In analogy to single polaron problem, we have $\hat{a}_{\mathbf{k}}\psi_{\rm X}=0$, where $\psi_{\rm X}=0$ is the exciton wave function. Then for the variational energy functional one has $E_{\rm X}=\langle\psi_{\rm X}|\hat{H}_{\rm X}|\psi_{\rm X}\rangle=\langle\psi_{\rm X}|\hat{H}_{\rm X0}|\psi_{\rm X}\rangle$. The ground state exciton wave function ($s$-orbital) assumed to be real, which results in the following relation
$\langle\psi_{\rm X}|\hat{p}\cdot\mathbf{j}+\mathbf{j}\cdot\hat{p}|\psi_{\rm X}\rangle=-i\hbar\int\psi_{\rm X}\left(2\mathbf{j}\cdot\nabla\psi_{\rm X}+(\nabla\mathbf{j})\psi_{\rm X}\right){\rm d}^{2}r=-i\hbar\int\left(2\psi_{\rm X}\mathbf{j}\cdot\nabla\psi_{\rm X}+(\nabla\mathbf{j})\psi^{2}_{\rm X}\right){\rm d}^{2}r$
$=-i\hbar\int\left(2\psi_{\rm X}\mathbf{j}\cdot\nabla\psi_{\rm X}+\nabla(\mathbf{j}\psi^{2}_{\rm X})-\mathbf{j}\cdot\nabla(\psi^{2}_{\rm X})\right){\rm d}^{2}r=-i\hbar\int\left(2\psi_{\rm X}\mathbf{j}\cdot\nabla\psi_{\rm X}-\mathbf{j}\cdot 2\psi_{\rm X}\nabla\psi_{\rm X}\right){\rm d}^{2}r=0$ (S122)
Setting the center-of-mass momentum $\mathbf{Q}=0$, we then get
$E_{\rm X}$ $=\langle\psi_{\rm X}|V_{\rm KR}(r,r_{\infty})+\frac{\hat{\mathbf{p}}^{2}}{2\mu}+\frac{\hbar^{2}}{2\mu}\sum_{k}\nabla F^{*}_{\mathbf{k}}\cdot\nabla F_{\mathbf{k}}+\sum_{k}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)|F_{\mathbf{k}}|^{2}-\sum_{k}\left(V_{k}\rho_{\mathbf{k}}F_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}F_{\mathbf{k}}^{*}\right)|\psi_{\rm X}\rangle,$ (S123)
where we retain the terms up to $|F|^{2}$. We now proceed with an approximate solution, neglecting the mixing of displacement $F_{\mathbf{k}}$ and exciton wave function $\psi_{\rm X}$. Then one can minimize the following functional:
$W_{\rm X}$ $=\frac{\hbar^{2}}{2\mu}\sum_{k}\nabla F^{*}_{\mathbf{k}}\cdot\nabla F_{\mathbf{k}}+\sum_{k}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)|F_{\mathbf{k}}|^{2}-\sum_{k}\left(V_{k}\rho_{\mathbf{k}}F_{\mathbf{k}}+V_{k}^{*}\rho_{\mathbf{k}}^{*}F_{\mathbf{k}}^{*}\right),$ (S124)
$\frac{\delta W_{\rm X}}{\delta F_{\mathbf{k}}}=\frac{\delta W_{\rm X}}{\delta F^{*}_{\mathbf{k}}}=0.$ (S125)
Evaluating, we get
$-\frac{\hbar^{2}}{2\mu}\Delta F_{\mathbf{k}}+\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)F_{\mathbf{k}}-V_{k}^{*}\rho_{\mathbf{k}}^{*}=0.$ (S126)
We employ a variational anzats
$F_{\mathbf{k}}=A_{\mathbf{k}}e^{-im_{h}\mathbf{k}\cdot\mathbf{r}/M}-B_{\mathbf{k}}e^{im_{e}\mathbf{k}\cdot\mathbf{r}/M}$ (S127)
yielding in
$\nabla F_{\mathbf{k}}$ $=-i\mathbf{k}\frac{m_{h}}{M}A_{\mathbf{k}}e^{-im_{h}\mathbf{k}\cdot\mathbf{r}/M}-i\mathbf{k}\frac{m_{e}}{M}B_{\mathbf{k}}e^{im_{e}\mathbf{k}\cdot\mathbf{r}/M},$ (S128)
$\Delta F_{\mathbf{k}}$ $=-k^{2}\frac{m_{h}^{2}}{M^{2}}A_{\mathbf{k}}e^{-im_{h}\mathbf{k}\cdot\mathbf{r}/M}+k^{2}\frac{m_{e}^{2}}{M^{2}}B_{\mathbf{k}}e^{im_{e}\mathbf{k}\cdot\mathbf{r}/M}.$ (S129)
Plugging in, we get
$\left(\frac{\hbar^{2}}{2\mu}k^{2}\frac{m_{h}^{2}}{M^{2}}+\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)A_{\mathbf{k}}e^{-im_{h}\mathbf{k}\cdot\mathbf{r}/M}-\left(\frac{\hbar^{2}}{2\mu}k^{2}\frac{m_{e}^{2}}{M^{2}}+\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)B_{\mathbf{k}}e^{im_{e}\mathbf{k}\cdot\mathbf{r}/M}=V_{k}^{*}\left(e^{-im_{h}\mathbf{k}\cdot\mathbf{r}/M}-e^{im_{e}\mathbf{k}\cdot\mathbf{r}/M}\right)$ (S130)
One has
$\frac{\hbar^{2}}{2\mu}k^{2}\frac{m_{h}^{2}}{M^{2}}+\frac{\hbar^{2}k^{2}}{2M}$ $=\frac{\hbar^{2}k^{2}}{2M}\left(\frac{m_{h}^{2}M}{m_{e}m_{h}M}+1\right)=\frac{\hbar^{2}k^{2}}{2M}\frac{m_{h}+m_{e}}{m_{e}}=\frac{\hbar^{2}k^{2}}{2m_{e}},$ (S131)
$\frac{\hbar^{2}}{2\mu}k^{2}\frac{m_{e}^{2}}{M^{2}}+\frac{\hbar^{2}k^{2}}{2M}$ $=\frac{\hbar^{2}k^{2}}{2M}\left(\frac{m_{e}^{2}M}{m_{e}m_{h}M}+1\right)=\frac{\hbar^{2}k^{2}}{2M}\frac{m_{e}+m_{h}}{m_{h}}=\frac{\hbar^{2}k^{2}}{2m_{h}},$ (S132)
so that
$A_{\mathbf{k}}$ $=\frac{V_{k}^{*}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}},\right)}$ (S133)
$B_{\mathbf{k}}$ $=\frac{V_{k}^{*}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)}.$ (S134)
Plugging into the energy functional, we reach at
$W_{\rm X}$ $=\sum_{k}\frac{\hbar^{2}k^{2}}{2M}|V_{k}|^{2}\left(\frac{m_{h}}{m_{e}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}+\frac{m_{e}}{m_{h}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)^{2}}+\frac{2\cos(\mathbf{k}\cdot\mathbf{r})}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)}\right)$
$+\sum_{k}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2M}\right)|V_{k}|^{2}\left(\frac{1}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}+\frac{1}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)^{2}}-\frac{2\cos(\mathbf{k}\cdot\mathbf{r})}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)}\right)$
$-\sum_{k}2|V_{k}|^{2}\left(\frac{1-\cos(\mathbf{k}\cdot\mathbf{r})}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{1-\cos(\mathbf{k}\cdot\mathbf{r})}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}\right)$
$=-\sum_{k}|V_{k}|^{2}\left(\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}\right)$
$+\sum_{k}2|V_{k}|^{2}\cos(\mathbf{k}\cdot\mathbf{r})\left(-\frac{\hbar\omega_{{\rm l},k}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)}+\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}\right)$
$=-\sum_{k}|V_{k}|^{2}\left(\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}\right)$
$+\sum_{k}|V_{k}|^{2}\cos(\mathbf{k}\cdot\mathbf{r})\left(\frac{m_{e}+m_{h}}{m_{e}-m_{h}}\left(\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}-\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}\right)+\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}\right)$
$=-\sum_{k}|V_{k}|^{2}\left(\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{1}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}\right)+\sum_{k}2|V_{k}|^{2}\frac{\cos(\mathbf{k}\cdot\mathbf{r})}{m_{h}-m_{e}}\left(\frac{m_{h}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}-\frac{m_{e}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\right).$ (S135)
The first term corresponds to single polaron binding energies of electron and hole. The second term results in the re-scaled Coulomb interaction, and is evaluated as
$\frac{\Sigma}{(2\pi)^{2}}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{\Sigma(4\varepsilon_{0})}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{2\cos(kr\cos\varphi)}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\left(\frac{m_{h}}{\hbar\omega_{\rm t}\sqrt{\frac{\varepsilon+r_{0}k}{\varepsilon+r_{\infty}k}}+\frac{\hbar^{2}k^{2}}{2m_{h}}}-\frac{m_{e}}{\hbar\omega_{\rm t}\sqrt{\frac{\varepsilon+r_{0}k}{\varepsilon+r_{\infty}k}}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\right)\frac{k{\rm d}k{\rm d}\varphi}{m_{h}-m_{e}}$
$=\frac{1}{(4\pi)}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{4\pi\varepsilon_{0}}\int\limits_{0}^{\infty}\frac{4\pi J_{0}(kr)}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\frac{\sqrt{\varepsilon+r_{\infty}k}}{\hbar\omega_{\rm t}\sqrt{\varepsilon+r_{0}k}}\left(\frac{m_{h}}{1+\frac{\hbar k^{2}}{2m_{h}\omega_{\rm t}}\sqrt{\frac{\varepsilon+r_{\infty}k}{\varepsilon+r_{0}k}}}-\frac{m_{e}}{1+\frac{\hbar k^{2}}{2m_{e}\omega_{\rm t}}\sqrt{\frac{\varepsilon+r_{\infty}k}{\varepsilon+r_{0}k}}}\right)\frac{k{\rm d}k}{m_{h}-m_{e}}$
$=\frac{e^{2}(r_{0}-r_{\infty})}{4\pi\varepsilon_{0}}\int\limits_{0}^{\infty}\frac{J_{0}(kr)}{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)}\left(\frac{m_{h}}{1+r_{h,{\rm t}}^{2}k^{2}\sqrt{\frac{\varepsilon+r_{\infty}k}{\varepsilon+r_{0}k}}}-\frac{m_{e}}{1+r_{e,{\rm t}}^{2}k^{2}\sqrt{\frac{\varepsilon+r_{\infty}k}{\varepsilon+r_{0}k}}}\right)\frac{k{\rm d}k}{m_{h}-m_{e}}$
$=\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\int\limits_{0}^{\infty}\frac{2J_{0}(\tilde{k}\varepsilon r/r_{\infty})}{(1+\sigma_{0}\tilde{k})(1+\tilde{k})}\left(\frac{m_{h}}{1+\sigma_{h,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}-\frac{m_{e}}{1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}\right)\frac{\tilde{k}{\rm d}\tilde{k}}{m_{h}-m_{e}}$ (S136)
Therefore, the effective Coulomb interaction is
$V_{\rm eff}(r)$ $=V_{\rm KR}(r,r_{\infty})+\sum_{k}2|V_{k}|^{2}\frac{\cos(\mathbf{k}\cdot\mathbf{r})}{\Delta m}\left(\frac{m_{h}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}}-\frac{m_{e}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\right)$
$=V_{\rm KR}(r,r_{\infty})+\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\int\limits_{0}^{\infty}\frac{2J_{0}(\tilde{k}\varepsilon r/r_{\infty})}{(1+\sigma_{0}\tilde{k})(1+\tilde{k})}\left(\frac{m_{h}}{1+\sigma_{h,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}-\frac{m_{e}}{1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}\right)\frac{\tilde{k}{\rm d}\tilde{k}}{\Delta m},$ (S137)
where $\Delta m=m_{h}-m_{e}$. We then apply an approximation $\sqrt{(1+\tilde{k})/(1+\sigma_{0}\tilde{k})}\approx 1/\sqrt{\sigma_{0}}$, and factorize the integrand:
$\frac{2\tilde{k}}{(1+\sigma_{0}\tilde{k})(1+\tilde{k})}\frac{1}{1+\epsilon_{i}\tilde{k}^{2}}=\frac{2}{(k+1)(\sigma_{0}-1)(\epsilon_{i}+1)}-\frac{2\sigma_{0}^{2}}{(\sigma_{0}-1)(k\sigma_{0}+1)\left(\sigma_{0}^{2}+\epsilon_{i}\right)}+\frac{2\left(k\epsilon_{i}^{2}-k\sigma_{0}\epsilon_{i}+\sigma_{0}\epsilon_{i}+\epsilon_{i}\right)}{(\epsilon_{i}+1)\left(k^{2}\epsilon_{i}+1\right)\left(\sigma_{0}^{2}+\epsilon_{i}\right)}$ (S138)
where $\epsilon_{i}=\sigma_{i,{\rm t}}^{2}/\sqrt{\sigma_{0}}$, $i=e,h$. Evaluating the terms one by one, we get
$\int\limits_{0}^{\infty}\frac{2J_{0}(\tilde{k}\varepsilon r/r_{\infty})}{(k+1)(\sigma_{0}-1)(\epsilon_{i}+1)}{\rm d}\tilde{k}$ $=\frac{\pi}{(\sigma_{0}-1)(\epsilon_{i}+1)}\left[H_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)\right],$ (S139)
$\int\limits_{0}^{\infty}\frac{2\sigma_{0}^{2}J_{0}(\tilde{k}\varepsilon r/r_{\infty})}{(\sigma_{0}-1)(k\sigma_{0}+1)\left(\sigma_{0}^{2}+\epsilon_{i}\right)}{\rm d}\tilde{k}$ $=\frac{\pi\sigma_{0}}{(\sigma_{0}-1)(\epsilon_{i}+\sigma_{0}^{2})}\left[H_{0}\left(\frac{\varepsilon r}{r_{\infty}\sigma_{0}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{\infty}\sigma_{0}}\right)\right],$ (S140)
$\int\limits_{0}^{\infty}J_{0}(\tilde{k}\varepsilon r/r_{\infty})\frac{2\epsilon_{i}\left(k\epsilon_{i}-k\sigma_{0}+\sigma_{0}+1\right)}{(\epsilon_{i}+1)\left(k^{2}\epsilon_{i}+1\right)\left(\sigma_{0}^{2}+\epsilon_{i}\right)}{\rm d}\tilde{k}$ $=\frac{1}{(1+\epsilon_{i})(\epsilon_{i}+\sigma_{0}^{2})}\left(\pi(1+\sigma_{0})\sqrt{\epsilon_{i}}\left[I_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{i}}}\right)-L_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{i}}}\right)\right]\right.$
$\left.+2(\epsilon_{i}-\sigma_{0})K_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{i}}}\right)\right),$ (S141)
where $I_{0}$, $K_{0}$, are modified Bessel functions, and $L_{0}$ is the Struve function. Hence, the overall potential is
$V_{\rm eff}(r)$ $=V_{\rm KR}(r,r_{\infty})+\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\left(i_{1}+i_{2}+i_{3}\right),$ (S142)
where
$i_{1}$ $=\frac{\pi}{2}\left[\frac{m_{h}}{\Delta_{m}(\epsilon_{h}+1)}-\frac{m_{e}}{\Delta_{m}(\epsilon_{e}+1)}\right]\left[H_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)\right],$ (S143)
$i_{2}$ $=-\frac{\pi}{2}\left[\frac{\sigma_{0}m_{h}}{\Delta_{m}(\epsilon_{h}+\sigma_{0}^{2})}-\frac{\sigma_{0}m_{e}}{\Delta_{m}(\epsilon_{e}+\sigma_{0}^{2})}\right]\left[H_{0}\left(\frac{\varepsilon r}{r_{0}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{0}}\right)\right],$ (S144)
$i_{3}$ $=\frac{\sigma_{0}-1}{2}\frac{m_{h}}{\Delta m}\frac{1}{(1+\epsilon_{h})(\epsilon_{h}+\sigma_{0}^{2})}\left(\pi(1+\sigma_{0})\sqrt{\epsilon_{h}}\left[I_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{h}}}\right)-L_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{h}}}\right)\right]+2(\epsilon_{h}-\sigma_{0})K_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{h}}}\right)\right)$
$-\frac{\sigma_{0}-1}{2}\frac{m_{e}}{\Delta m}\frac{1}{(1+\epsilon_{e})(\epsilon_{e}+\sigma_{0}^{2})}\left(\pi(1+\sigma_{0})\sqrt{\epsilon_{e}}\left[I_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{e}}}\right)-L_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{e}}}\right)\right]+2(\epsilon_{e}-\sigma_{0})K_{0}\left(\frac{\varepsilon r}{r_{\infty}\sqrt{\epsilon_{e}}}\right)\right)$ (S145)
We note that $\epsilon_{i}\ll 1$, and expand to first order, yielding in
$i_{1}$ $\approx\frac{\pi}{2}\frac{m_{h}-m_{e}-m_{h}\epsilon_{h}+m_{e}\epsilon_{e}}{\Delta_{m}}\left[H_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)\right]=\frac{\pi}{2}\left[H_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{\infty}}\right)\right],$ (S146)
$i_{2}$ $\approx-\frac{\pi}{2}\frac{m_{h}-m_{e}-(m_{h}\epsilon_{h}-m_{e}\epsilon_{e})/\sigma_{0}^{2}}{\Delta_{m}\sigma_{0}}\left[H_{0}\left(\frac{\varepsilon r}{r_{0}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{0}}\right)\right]=-\frac{\pi}{2\sigma_{0}}\left[H_{0}\left(\frac{\varepsilon r}{r_{0}}\right)-Y_{0}\left(\frac{\varepsilon r}{r_{0}}\right)\right],$ (S147)
where we use
$m_{h}\epsilon_{h}-m_{e}\epsilon_{e}=(m_{h}\sigma_{h,{\rm t}}^{2}-m_{e}\sigma_{e,{\rm t}}^{2})/\sqrt{\sigma_{0}}=0.$ (S148)
One can now note the cancellation
$V_{\rm KR}(r,r_{\infty})+\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}i_{1}=0,$ (S149)
which eventually results in the following effective potential
$V_{\rm eff}(r)$ $\approx V_{\rm KR}(r,r_{0})-\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\left[\frac{m_{h}}{\Delta m}\frac{\pi(1+\sigma_{0})\tilde{\sigma}_{h,{\rm t}}\left(I_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)-L_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)\right)+2(\tilde{\sigma}_{h,{\rm t}}^{2}-\sigma_{0})K_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)}{(1+\tilde{\sigma}_{h,{\rm t}}^{2})(\sigma_{0}^{2}+\tilde{\sigma}_{h,{\rm t}}^{2})}\right.$
$\hskip 85.0pt\left.-\frac{m_{e}}{\Delta m}\frac{\pi(1+\sigma_{0})\tilde{\sigma}_{e,{\rm t}}\left(I_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{e,{\rm t}}r_{\infty}}\right)-L_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{e,{\rm t}}r_{\infty}}\right)\right)+2(\tilde{\sigma}_{e,{\rm t}}^{2}-\sigma_{0})K_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{e,{\rm t}}r_{\infty}}\right)}{(1+\tilde{\sigma}_{e,{\rm t}}^{2})(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})}\right],$ (S150)
where we recall $\tilde{\sigma}_{i,{\rm t}}=\sigma_{i,{\rm t}}/\sigma_{0}^{1/4}$, $i=e,h$. We can further simplify to the form
$V_{\rm eff}(r)$ $\approx V_{\rm KR}(r,r_{0})-E_{0}\left[\frac{m_{h}}{\Delta m}h_{\tilde{\sigma}_{h,{\rm t}}}(\varepsilon r/r_{\infty})-\frac{m_{e}}{\Delta m}h_{\tilde{\sigma}_{e,{\rm t}}}(\varepsilon r/r_{\infty})\right],$ (S151)
where
$h_{\nu}(x)=\dfrac{\sigma_{0}-1}{2(1+\nu^{2})(\sigma_{0}^{2}+\nu^{2})}\left[(1+\sigma_{0})\alpha\pi\left[I_{0}\left(x/\nu\right)-L_{0}\left(x/\nu\right)\right]+2(\nu^{2}-\sigma_{0})K_{0}\left(x/\nu\right)\right].$ (S152)